### v2

# The synthetic instrument: From sparse association to sparse causation

Abstract
In many observational studies, researchers are often interested in studying the effects of multiple exposures on a single outcome. Standard approaches for high-dimensional data, such as the lasso, assume that the associations between the exposures and the outcome are sparse. However, these methods do not estimate causal effects in the presence of unmeasured confounding. In this paper, we consider an alternative approach that assumes the causal effects under consideration are sparse. We show that with sparse causation, causal effects are identifiable even with unmeasured confounding. At the core of our proposal is a novel device called the synthetic instrument, which, in contrast to standard instrumental variables, can be constructed using the observed exposures directly. We demonstrate that under the assumption of sparse causation, the problem of causal effect estimation can be formulated as an $\ell_{0}$-penalization problem and can thus be solved efficiently using off-the-shelf software. Simulations show that our approach outperforms state-of-the-art methods in both low-dimensional and high-dimensional settings. We further illustrate our method using a mouse obesity dataset.
Dingke Tang1, Dehan Kong2, and Linbo Wang2***Address for correspondence: Linbo Wang, Department of Statistical Sciences, University of Toronto, 700 University Avenue, 9th Floor, Toronto, ON, Canada, M5G 1Z5
Email: linbo.wang@utoronto.ca
1Department of Statistics and Data Science, Washington University in St. Louis, St. Louis, Missouri, U.S.A.
2 Department of Statistical Sciences, University of Toronto, Toronto, Ontario, Canada
Keywords: Causal inference; Multivariate analysis; Unmeasured confounding.
1 Introduction
Sparsity is a common assumption in modern statistical learning literature, as it allows for variable selection in underlying models and enhances the interpretability of parameter estimates. For example, the lasso method (Tibshirani, 1996) assumes that the underlying associations between a single outcome and potentially high-dimensional predictors are sparse. In other words, only a small number of predictors have non-zero associations with the outcome. Hastie et al. (2009) summarize the philosophy behind these methods as the “bet on sparsity” principle: Use a procedure that performs well in sparse problems, as no procedure performs well in dense problems. Methods like the lasso perform well when the associations between the outcome and predictors are sparse and have gained significant popularity in recent decades.
The “bet on sparsity” principle does not restrict the type of “sparse problems” that can be considered. Instead of focusing solely on sparse associations, there is a growing body of literature that emphasizes sparse causation, where only a fraction of the exposures have non-zero causal effects on the outcomes (e.g., Spirtes and Glymour, 1991; Claassen et al., 2013; Wang et al., 2017; Miao et al., 2023; Zhou et al., 2024). This assumption is arguably more interpretable and plausible in many data applications. For example, suppose one is interested in the relationships between the expression of multiple genes and a phenotype of interest, such as lung cancer. Current biological knowledge suggests that not all genes in the human genome may affect lung cancer (e.g., Kanwal et al., 2017). However, this does not imply the assumption of sparse association, as unmeasured confounding often leads to spurious correlations between the exposures and the outcome.
For example, consider a linear structural model (Pearl, 2013) with a $p$-dimensional exposure $X=(X_{1},\ldots,X_{p})^{\mathrm{\scriptscriptstyle T}}$, an outcome $Y$, and a $q$-dimensional latent variable $U$:
$X$ $=\Lambda U+\epsilon_{x},$ (1)
$Y$ $=X^{\mathrm{\scriptscriptstyle T}}\beta+U^{\mathrm{\scriptscriptstyle T}}\gamma+\epsilon_{y},$ (2)
where $\Lambda\in\mathbb{R}^{p\times q}$, $\beta\in\mathbb{R}^{p}$, and $\gamma\in\mathbb{R}^{q}$ are coefficients, and $\epsilon_{x}=(\epsilon_{1},\ldots,\epsilon_{p})^{\mathrm{\scriptscriptstyle T}}$, $\epsilon_{y}$, and $U$ are uncorrelated with each other. Under this model, the spurious correlations due to unmeasured confounding, characterized by $\text{Cov}(X)^{-1}\Lambda\gamma$, are typically dense. As a result, the association between $X$ and $Y$ is not sparse, even if the causal effect $\beta$ is sparse.
Identification and estimation of the causal parameter $\beta$ are non-trivial due to the presence of unmeasured confounding by $U$. Our contributions in this paper are twofold. First, under an additional plurality condition, we establish that the parameter $\beta$ in model (3) is identifiable if and only if $||\beta||_{0}<p-q$. Our assumption on the sparsity level is both necessary and sufficient, representing a significant improvement over assumptions previously introduced in the literature; see Section 1.1 for details. Remarkably, in contrast to many other identification assumptions in causal inference, it is possible to consistently test the sparsity assumption from data.
Second, we develop a synthetic two-stage regularized regression approach for estimating $\beta$, with a first-stage ordinary least squares and a second-stage $\ell_{0}$-penalized regression. The key technique behind our results is a novel device we call the synthetic instrument. Unlike standard instrumental variables, the synthetic instrument is constructed from a subset of exposures, allowing us to identify the causal effects without resorting to exogenous variables. Our procedure comes with lasso-type theoretical guarantees in both low- and high-dimensional settings.
1.1 Related works
Our proposal is related to recent work dealing with multivariate hidden confounding. Ćevid et al. (2020) and Guo et al. (2022) propose a spectral deconfounding method for estimating $\beta$ under a high-dimensional model. Their method assumes a dense confounding model, which is only possible under a high-dimensional regime where $p$ tends to infinity with the sample size and the magnitude of spurious associations tends to zero. Bing et al. (2022) consider a more general setup than ours, in which they also allow the outcome to be multivariate. However, they aim to identify the projection of $\beta$ onto a related space rather than the causal parameter $\beta$ itself. Chandrasekaran et al. (2010) study a related problem under the assumption that $(X,Y)$ are normally distributed. Under this assumption, they not only identify the effect of $X$ on $Y$ but also recover the covariance among components of $X$ conditional on $U$.
The estimation problem of $\beta$ can be framed within the context of causal inference with unmeasured confounding. Currently, the most popular approach in practice is the instrumental variable (IV) framework, which uses information from an exogenous variable known as an IV to identify causal effects. (e.g., Angrist et al., 1996; Wang and Tchetgen Tchetgen, 2018). Another approach that has gained attention recently is the proximal causal inference framework (Tchetgen Tchetgen et al., 2024), which uses information from ancillary variables known as negative control exposures and outcomes to remove bias due to unmeasured confounding. Compared with these frameworks, our approach does not rely on the collection of additional ancillary variables, which can be challenging in many practical settings. Instead, we rely on the availability of multiple exposures and the sparsity assumption for identification and estimation.
Recently, there has been a strand of literature that seeks to identify the causal effects of multiple exposures. Wang and Blei (2019) popularized this setting by proposing the so-called deconfounder method, which first obtains an estimate $\widehat{U}$ for the unmeasured confounder and then adjusts for $\widehat{U}$ using standard regression methods. However, it has been pointed out that under this setting, without further assumptions, the causal effect $\beta$ is not identifiable (D’Amour, 2019; Ogburn et al., 2020). Kong et al. (2022) show that under model (1) and a binary choice model for the outcome with a non-probit link, the causal effects are identifiable. Their identification results, however, apply only to binary outcomes and do not lead to straightforward estimation procedures. Miao et al. (2023) consider a similar setting to (1) and (3), showing that the causal effect is identifiable if $\|\beta\|_{0}\leq(p-q)/2$. Their sparsity constraint is significantly stronger than ours, especially in cases where the number of exposures is large relative to the number of latent confounders. Miao et al. (2023) also develop a robust linear regression-based estimator for estimating $\beta$. In contrast to our estimator, their estimator is only consistent in the low-dimensional regime where $p$ is fixed and $\|\beta\|_{0}\leq p/2-q+1$. Furthermore, their estimator for $\beta$ is not sparse and, therefore, cannot be used for selecting treatments with non-zero effects.
Our results also connect to recent literature on multiply robust causal identification (e.g., Sun et al., 2023), as we show identification in the union of many causal models. This contrasts with the rich literature on multiply robust estimators under the same causal model (e.g., Wang and Tchetgen Tchetgen, 2018) and improved doubly robust estimators that are consistent under multiple working models for two components of the likelihood (e.g., Han and Wang, 2013).
1.2 Outline of this paper
The rest of this article is organized as follows. In Section 2, we introduce the setup and background. In Section 3, we introduce our identification strategy using the synthetic instrument method. In Section 4, we present our estimation procedure and provide theoretical justifications. We also discuss extensions to non-linear outcome models. Simulation studies in Section 5 compare our proposal with several state-of-the-art methods in terms of their finite-sample performance. In Section 6, we apply our method to mouse obesity data. We conclude with a brief discussion in Section 7.
The proposed method is implemented in an R package available at https://github.com/dingketang/syntheticIV.
2 Framework, notation, and identifiability
2.1 The model
We assume that we observe $n$ independent samples from the joint distribution of $(X,Y)$. Consider structural model (1) and
$Y$ $=X^{\mathrm{\scriptscriptstyle T}}\beta+g(U)+\epsilon_{y}.$ (3)
Here, $g(U):\mathbb{R}^{q}\to\mathbb{R}$ is a measurable function encoding the effects of unmeasured confounders $U$ on the outcome $Y$. We do not assume knowledge of the functional form of $g(U)$ because $U$ is unmeasured, making it implausible to specify the exact form of $g(\cdot)$.
We start with a linear outcome model where the treatment effect is linear in $X$. In Section 4.3, we will consider a nonlinear treatment effect model, where the relationship between treatment $X$ and outcome $Y$ is represented by a potentially nonlinear function $f(X;\beta)$.
We consider both low- and high-dimensional settings where $p$ may be smaller or larger than the sample size $n$. Assume $\dot{\beta}$ is the true value of $\beta$ in model (3). Without loss of generality, we assume all the variables in (1) and (3) are centered, $\text{Cov}(U)=I_{q}$, and $\mathbb{E}(g(U))=0$.
We maintain the following conditions throughout the article.
  1. A1
    (Invertibility) Any $q\times q$ submatrix of $\text{Cov}^{-1}(X)\Lambda$ is invertible.
  2. A2
    $\Lambda$ is identifiable up to a rotation.
Condition A1 is a regularity condition commonly assumed in the literature (e.g., Miao et al., 2023, Theorem 3). However, it may be relaxed in our setting. For example, if some treatment effects are not confounded after normalization, such that certain rows of $\text{Cov}^{-1}(X)\Lambda$ are zero, then although Condition A1 is violated, the proposed method can still be used to identify and estimate the treatment effects. See Section LABEL:sec:discussA1 of the supplementary material for more details.
Condition A2 has been discussed extensively in the factor model literature. One classical result is Proposition 1 below, which is a direct corollary of Anderson and Rubin (1956, Theorem 5.1).
Proposition 1.
Under models (1), (3), and Condition A1, if $p\geq 2q+1$ and $D=\text{Cov}(\epsilon_{x})$ is a diagonal matrix, then $\Lambda$ is identifiable up to a rotation.
We note that the condition that $D$ is a diagonal matrix is a classical assumption in the factor analysis literature. However, Condition A2, and hence our algorithm, may still hold even if $D$ is not diagonal. For example, under the assumption that $D$ is sparse, the covariance structure $\text{Cov}(X)=D+\Lambda\Lambda^{\mathrm{\scriptscriptstyle T}}$ implies a sparse plus low-rank decomposition. This allows for the identification of the low-rank component $\Lambda\Lambda^{\mathrm{\scriptscriptstyle T}}$, as established in Chandrasekaran et al. (2011, Corollary 3), which leads to identifying $\Lambda$ up to a rotation. In another example, in the high-dimensional setting where $p\rightarrow\infty$, it is possible to identify $\Lambda^{*}\in\mathbb{R}^{p\times q}$, whose columns correspond to the top $q$ eigenvalues of $\text{Cov}(X)$. Under additional boundedness assumptions on the correlation matrix $D$ and the coefficient matrix $\Lambda$, one can show that there exists a matrix $O\in\mathbb{R}^{q\times q}$ such that the $\ell_{2}$ norm between each column of $\Lambda O$ and $\Lambda^{*}$ converges to zero as $p$ tends to infinity. See Fan et al. (2013, Proposition 2.2, Theorem 3.3), Bai (2003, Theorem 2), and Shen et al. (2016, Theorem 1) for more details.
2.2 Identifiability of the causal effect $\beta$
In this section, we discuss the identifiability of the causal parameter $\beta$ in (3). We illustrate the key ideas using the specific example where $p=3$ and $q=1$ in models (1) and (3). Figure 1 provides graphical illustrations.
First, note that without additional assumptions, $\beta$ is generally not identifiable due to unmeasured confounding by $U$. To see this, observe that under models (1) and (3), we have
$\text{Cov}(X_{j},Y)=\beta_{1}\text{Cov}(X_{j},X_{1})+\beta_{2}\text{Cov}(X_{j},X_{2})+\beta_{3}\text{Cov}(X_{j},X_{3})+\gamma\Lambda_{j},\quad j=1,2,3,$ (4)
where $\Lambda_{j}$ is the $j$th element of $\Lambda\in\mathbb{R}^{p\times 1}$ and $\gamma=\mathbb{E}(Ug(U))$. Since there are three equations in (4) but four unknown parameters $\beta_{1},\beta_{2},\beta_{3},\gamma$, the causal parameters $\beta$ are not identifiable from these equations.
One possible approach to identifying $\beta$ is to assume prior knowledge about certain elements of $\beta$. For instance, in Figure 1(b), it is assumed that $\beta_{2}=0$, meaning that $X_{2}$ has no causal effect on the outcome $Y$. In this scenario, it is straightforward to see from (4) that under Conditions A1 and A2, $\beta_{1}$, $\beta_{3}$, and $|\gamma|$ are identifiable.
[…]
(a) No additional assumptions.
[…]
(b) Assume $\beta_{2}=0$.
[…]
(c) Assume $||\beta||_{0}\leq 1$.
Figure 1: Causal diagrams corresponding to models (1) and (3): $p=3,q=1.$
In practice, however, it is often difficult to know which exposures have zero causal effects a priori. In this paper, we instead consider the following sparsity assumption; see Figure 1c for an illustration.
  1. A3
    (Sparsity) $\|\dot{\beta}\|_{0}\leq p-q-1$, where $\dot{\beta}$ refers to the true value of $\beta$.
Remark 1.
Condition A3 is significantly less restrictive compared to similar assumptions in the existing literature used for causal effect identification in this context. For example, Miao et al. (2023) assumed $\|\dot{\beta}\|_{0}\leq(p-q)/2$.
2.3 Instrumental variable
The method of instrumental variables (IV) is a widely used approach for estimating causal relationships when unmeasured confounders exist between the exposure $X$ and the outcome $Y$. Suppose we have an exogenous variable $Z$. For simplicity, assume that the relationships among the random variables are linear and follow the structure equation models:
$\begin{split}Y&=\beta X+\gamma U+\pi Z+\epsilon_{y},\\ X&=\alpha_{z}Z+\Lambda U+\epsilon_{x}.\end{split}$
For $Z$ to be a valid instrumental variable, the following assumptions are commonly made (e.g., Wang and Tchetgen Tchetgen, 2018): $\pi=0$ (exclusion restriction), $\alpha_{z}\neq 0$ (instrumental relevance), and $\text{Cov}(U,Z)=0$ (unconfoundedness). Under these assumptions, one can consistently estimate $\beta$ via a two-stage least squares estimator: first, obtain the predicted exposure $\widehat{\mathbb{E}}(X\mid Z)$ by linearly regressing $X$ on $Z$, and then regress $Y$ on $\widehat{\mathbb{E}}(X\mid Z)$ to obtain an estimate of $\beta$; here $\mathbb{E}(X\mid Z)$ refers to the conditional expectation of $X$ given $Z$.
3 Identifying causal effects via the synthetic instrument
3.1 A new identification approach via voting
We now present a new identification strategy for $\beta$ under the sparsity condition A3. Consider the scenario depicted in Figure 1(c), where $\dot{\beta}_{1}=\dot{\beta}_{2}=0$ but $\dot{\beta}_{3}\neq 0$. We assume that this information is unavailable to the analyst. Instead, the analyst relies on Condition A3, assuming that $\|\dot{\beta}\|_{0}\leq 1$.
To explain the identification strategy, it is helpful to consider a voting analogy; see also Zhou et al. (2014) and Guo et al. (2018) for similar approaches in different contexts. Suppose the analyst consults three experts, and expert $j$ hypothesizes that $\beta_{j}=0$. Based on this hypothesis, one can identify the other elements in $\beta$ using the approach described in Section 2.2. Specifically, for $j=1,2,3,$ let $\widetilde{\beta}^{(j)}$ (and $|\widetilde{\gamma}^{(j)}|$) solve (4) assuming $\beta_{j}=0$. Table 1 summarizes these solutions. Note that the hypotheses by experts 1 and 2 are both correct, so we have $\widetilde{\beta}^{(1)}=\widetilde{\beta}^{(2)}=\beta$ under Conditions A1A2. On the other hand, the hypothesis postulated by expert 3 is incorrect. Therefore, in general, $\widetilde{\beta}^{(3)}\neq\beta$. To decide among these three experts, we compare the solutions $\widetilde{\beta}^{(j)}$ and find their mode $\beta_{\text{mode}}=\mathop{\arg\max}\limits_{\beta\in\mathbb{R}^{3}}|\{j:\widetilde{\beta}^{(j)}=\beta\}|$; here $|\mathcal{S}|$ denotes the cardinality of a set $\mathcal{S}$. One can easily see from Table 1 that $\beta_{\text{mode}}=\dot{\beta}.$
Table 1: A voting analogy of our identification approach for $\beta$. Note $\widetilde{\beta}^{(j)}=\left(\widetilde{\beta}^{(j)}_{1},\widetilde{\beta}^{(j)}_{2},\widetilde{\beta}^{(j)}_{3}\right)$ denotes the solution to equation (4) under the hypothesis that $\beta_{j}=0$
Expert index $j$ Expert hypothesis Solution to the identification equation (4)
$\widetilde{\beta}_{1}^{(j)}$ $\widetilde{\beta}_{2}^{(j)}$ $\widetilde{\beta}_{3}^{(j)}$
$j=1$ $\beta_{1}=0$ 0 0 $\dot{\beta}_{3}$
$j=2$ $\beta_{2}=0$ 0 0 $\dot{\beta}_{3}$
$j=3$ $\beta_{3}=0$ non-zero non-zero 0
3.2 The synthetic instrument
On the surface, one may follow the identification strategy described in Section 3.1 to estimate $\beta$. However, in the general case where $q>1$, each expert would hypothesize that exactly $q$ elements of $\beta$ are zero. In total, there are $C_{p}^{q}$ different hypotheses. Several challenges arise when the data is moderate to high-dimensional, so that $p$ and $q$ are not small.
  1. (i)
    One needs to solve the empirical version of equation (4) $C_{p}^{q}$ times. This could be computationally expensive.
  2. (ii)
    Finding the mode of $C_{p}^{q}$ $p$-dimensional estimates is a non-trivial statistical problem.
To overcome these challenges, we introduce a new device, called the synthetic instrumental variable (SIV) method. As we shall see later, the SIV method has significant advantages in terms of both computational efficiency and identifiability for $\beta$.
Remark 2.
Other approaches that use the voting analogy for identification (e.g. Zhou et al., 2014; Guo et al., 2018) face the same challenges we present here. It is only due to the special structure of our problem that we are able to develop a method that bypasses the model selection step and addresses these challenges. See Sun et al. (2023) for another example in a different context.
In the following, we first introduce the SIV in the context of Figure 1(b), where it is assumed that $\beta_{2}=0$. Note from Figure 1(b) that the error term $\epsilon_{1}$ serves as an instrumental variable for estimating the effect parameter $\beta_{1}$. However, $\epsilon_{1}$ is not observable. Instead, note that (1) implies
$\begin{split}X_{1}&=\Lambda_{1}U+\epsilon_{1},\\ X_{2}&=\Lambda_{2}U+\epsilon_{2},\end{split}$ (5)
where $\Lambda_{1}$ and $\Lambda_{2}$ are identified up to the same sign flip so that $\Lambda_{1}/\Lambda_{2}$ is identifiable. Eliminating $U$ from (5), we get $X_{1}-\frac{\Lambda_{1}}{\Lambda_{2}}X_{2}=\epsilon_{1}-\frac{\Lambda_{1}}{\Lambda_{2}}\epsilon_{2}$, which only depends on the error terms $\epsilon_{1}$ and $\epsilon_{2}$. Since $\epsilon_{2}$ is also uncorrelated with $U$, it is not difficult to see from Figure 1(b) that $SIV_{1}^{(2)}=X_{1}-\frac{\Lambda_{1}}{\Lambda_{2}}X_{2}$ satisfies the conditions for an instrumental variable for identifying $\beta_{1}$ described in Section 2.3, hence the name synthetic instrument. In contrast to a standard instrumental variable, the synthetic instrument is directly constructed as a linear combination of the exposures, so there is no need to measure additional exogenous variables.
To identify $\beta_{3}$, one can similarly define $SIV_{1}^{(3)}=X_{3}-\frac{\Lambda_{1}}{\Lambda_{2}}X_{3}.$ Let $SIV^{(2)}=\left(SIV_{1}^{(2)},SIV_{3}^{(2)}\right)$. One can then obtain $(\beta_{1},\beta_{3})$ using the so-called synthetic two-stage least squares:
  1. 1.
    Fit a linear regression of $X=(X_{1},X_{2},X_{3})$ on $SIV^{(2)}=(SIV_{1}^{(2)},SIV_{3}^{(2)})$ and obtain $\widetilde{X}=E[X\mid SIV^{(2)}]$.
  2. 2.
    Fit a linear regression of $Y$ on $\widetilde{X}$, fixing $\beta_{2}=0$, and obtain the coefficients $\widetilde{\beta}_{1}$ and $\widetilde{\beta}_{3}$.
3.3 Voting with the synthetic instrument
Now consider applying the synthetic instrument to the case in Figure 1(c), where the analyst does not have prior information on which exposure has zero effect on the outcome. Instead, we assume the sparsity condition that $\|\beta\|_{0}\leq 1$.
By combining the voting procedure in Section 3.1 with the synthetic two-stage least squares method in Section 3.2, we arrive at Algorithm 1 for the estimation of $\beta$.
Algorithm 1 A naive voting procedure with synthetic two-stage least squares
  1. 1.
    For $j=1,2,3$, fit a linear regression of $X$ on $SIV^{(j)}$ and obtain $\widetilde{X}^{(j)}=E\left[X\mid SIV^{(j)}\right]$;
  2. 2.
    Fit a linear regression of $Y$ on $\widetilde{X}^{(j)}$, fixing $\beta_{j}=0$, and obtain the coefficients $\widetilde{\beta}^{(j)}$.
  3. 3.
    Find the mode among $\widetilde{\beta}^{(j)},j=1,2,3.$
On the surface, similar to the problems described at the beginning of Section 3.2, voting with the synthetic instrument still involves fitting three different regressions and comparing three vectors $\widetilde{\beta}^{(j)}.$ We now make two key observations regarding the properties of the synthetic instrument, which allows us to simplify Algorithm 1 into a two-stage regression procedure.
Observation 1.
Let $\Lambda=(\Lambda_{1},\Lambda_{2},\Lambda_{3})$. For $j=1,2,3$, $SIV^{(j)}\in\mathbb{R}^{2}$ span the same linear space $\{\lambda^{T}X:\lambda\in\Lambda^{\perp}\}.$ As a result, $\widetilde{X}^{(j)}=E\left[X\mid SIV^{(j)}\right]$ does not depend on the choice of $j$ so that one only needs to run Step 1 of Algorithm 1 once.
Observation 2.
From Table 1, we observe that $\|\widetilde{\beta}^{(1)}\|_{0}=\|\widetilde{\beta}^{(2)}\|_{0}=1$, while $\|\widetilde{\beta}^{(3)}\|_{0}=2$. Recall that the true value is $\dot{\beta}=\widetilde{\beta}^{(1)}=\widetilde{\beta}^{(2)}$. Instead of calculating $\widetilde{\beta}^{(j)},j=1,2,3$ separately for each $j$, Steps 2 and 3 in Algorithm 1 can be replaced with the following penalized regression:
$\beta^{SIV}=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{3}}\|\ Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta\|_{2}^{2}\text{\quad subject to }\|\beta\|_{0}\leq 1,$
where, due to Observation 1, $\widetilde{X}^{(1)}=\widetilde{X}^{(2)}=\widetilde{X}^{(3)}\equiv\widetilde{X}.$
With these observations, Algorithm 1 simplifies to a two-step regularized regression procedure.
3.4 Synthetic two-stage regularized regression
We now formally introduce the synthetic two-stage regularized regression for the general case. Motivated by Observation 1, we introduce the following definition for the synthetic instrument.
Definition 1 (Synthetic Instrument).
Define
$SIV=B_{\Lambda^{\perp}}^{\mathrm{\scriptscriptstyle T}}X\in\mathbb{R}^{p-q},$
where $B_{\Lambda^{\perp}}\in\mathbb{R}^{p\times(p-q)}$ is a semi-orthogonal matrix whose column space is orthogonal to the column space of $\Lambda\in\mathbb{R}^{p\times q}$.
The following proposition confirms that the $SIV$ are valid instruments.
Proposition 2.
Under models (1), (3), and Condition A2, the $SIV$ given by Definition 1 serve as valid instrumental variables for estimating the treatment effects of $X$ on $Y$.
To identify the causal parameter $\beta$ in the general case, we introduce the following plurality condition A4.
  1. A4
    (Plurality rule) Let $C^{*}$ be a subset of $\{1,2,\ldots,p\}$ with cardinality $q$, and suppose that $\dot{\beta}_{C^{*}}\neq 0$. The synthetic two-stage least squares coefficient obtained by assuming $\beta_{C^{*}}=0$ is given by $\widetilde{\beta}^{C^{*}}=\underset{\beta\in\mathbb{R}^{p}:\beta_{C^{*}}=0}{\operatorname*{arg\,min}}\,\mathbb{E}(Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta)^{2}$, where $\widetilde{X}=\mathbb{E}(X\mid SIV)$. The plurality rule assumes that $\max\limits_{\beta\in\mathbb{R}^{p}}|\{C^{*}:\widetilde{\beta}^{C^{*}}=\beta\}|\leq q$.
In Condition A4, each $C^{*}$ corresponds to an expert who makes the incorrect hypothesis that $\beta_{C^{*}}=0$. Let $s=\|\dot{\beta}\|_{0}$. In general, there are $C_{p-s}^{q}$ experts making correct hypotheses. If $s<p-q$, then there are at least $q+1$ experts making correct hypotheses. The plurality rule assumes that no more than $q$ incorrect hypotheses lead to the same synthetic two-stage least squares coefficient. This assumption is similar in spirit to the plurality assumption used in the invalid IV literature (e.g. Guo et al., 2018). We discussed this assumption in more detail in Section LABEL:sec:disA4 of the supplementary material.
In parallel to Observation 2, we have the following theorem.
Theorem 1 (Synthetic two-stage regularized regression).
Suppose that models (1), (3) and conditions A1, A2, and A4 hold.
  1. 1.
    If A3 holds, then $\dot{\beta}$ is identifiable via $\dot{\beta}=\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;\mathbb{E}(Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta)^{2},$ subject to $\|\beta\|_{0}<p-q$.
  2. 2.
    If A3 fails, then $\dot{\beta}$ is not identifiable, and for all $\widetilde{\beta}\in\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;\mathbb{E}(Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta)^{2}$, we have $\|\widetilde{\beta}\|_{0}\geq p-q$.
An important feature of Theorem 1 is that given $q$, it is possible to test the sparsity condition A3 from the observed data. In particular, it shows that under models (1), (3), conditions A1, A2, and the plurality rule A4, the following three statements are equivalent:
  1. (1)
    $\beta$ is identifiable;
  2. (2)
    Condition A3 holds;
  3. (3)
    The most sparse least-squares solution to the second-stage regression has an $\ell_{0}$-norm smaller than $p-q$, i.e.,
    $\min\limits_{\widetilde{\beta}\in\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;\mathbb{E}(Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta)^{2}}\|\widetilde{\beta}\|_{0}<p-q.$ (6)
It is worth noting that (6) can be checked from the observed data distribution, so that one may develop a consistent test for Condition A3 and identifiability of $\beta$ under models (1), (3) and conditions A1, A2, and A4. See Algorithm 2 below for more details.
4 Estimation via the synthetic two-stage regularized regression
4.1 Estimation
Let ${\bf X}\in\mathbb{R}^{n\times p}$ be the design matrix and ${\bf{Y}}\in\mathbb{R}^{n\times 1}$ denote the observed outcome. Theorem 1 suggests the following synthetic two-stage regularized regression for estimating $\beta$:
$\begin{split}\widehat{\bf X}&=\widehat{\mathbb{E}}({\bf{X}}\mid\widehat{SIV}),\\ \widehat{\beta}&=\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}{\|{\bf{Y}}-\widehat{{\bf X}}\beta\|_{2}^{2}}\quad\text{subject to }\|\beta\|_{0}\leq k,\end{split}$ (7)
where $k$ is a tuning parameter, $\widehat{\Lambda}$ is an estimator of the loading matrix, and $\widehat{SIV}=\bm{X}B_{\widehat{\Lambda}^{\perp}}.$
Several different estimators have been proposed for determining the number of latent factors in a factor model. In our simulations and data analysis, we will use the estimator developed by Onatski (2010) to obtain $\widehat{q}$, as it is applicable in both low- and high-dimensional settings. Similarly, numerous estimators have been developed for the loading matrix $\Lambda$. For low-dimensional settings where $p$ is fixed, we recommend using the maximum likelihood estimator for $\Lambda$ by optimizing the log-likelihood under the assumption of multivariate normality. For high-dimensional settings, we suggest estimating $\Lambda$ using the principal component analysis (PCA) method (Bai, 2003), which provides a consistent estimator for the loading matrix without assuming the covariance matrix $\text{Cov}(\epsilon_{X})$ to be diagonal. Finally, we use cross-validation to select the tuning parameter $k$. Algorithm 2 summarizes our estimation procedure.
Algorithm 2  The synthetic two-stage regularized regression
Input: ${\bf{X}}\in\mathbb{R}^{n\times p}$(centered), ${\bf{Y}}\in\mathbb{R}^{n\times 1}$
1:Obtain $\widehat{q}$ from $\bm{X}$ (e.g., Onatski, 2010).
2:if $n>p$ then obtain $\widehat{\Lambda}\in\mathbb{R}^{p\times q}$ via maximum likelihood estimation, assuming multivariate normality;
3:else let $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\geq\ldots\geq\widehat{\lambda}_{p}$ be the eigenvalues of ${\bf X}^{\mathrm{\scriptscriptstyle T}}{\bf X}/(n-1)$, and let $\widehat{\xi}_{1},\widehat{\xi}_{2},\ldots,\widehat{\xi}_{p}$ be the corresponding eigenvectors. Define $\widehat{\Lambda}=(\sqrt{\widehat{\lambda}_{1}}\widehat{\xi}_{1}\;\ldots\;\sqrt{\widehat{\lambda}_{q}}\widehat{\xi}_{q})$.
4:Let $B_{\widehat{\Lambda}^{\perp}}$ be a semi-orthogonal matrix whose columns are orthogonal to the columns of $\widehat{\Lambda}$. This can be obtained, for example, using the Null function from the MASS package in R.
5:Obtain $\widehat{SIV}=\bm{X}B_{\widehat{\Lambda}^{\perp}}$.
6:Obtain $\widehat{\bf X}=\widehat{\mathbb{E}}({\bf{X}}\mid\widehat{SIV})$ via ordinary least squares.
7:Obtain $\widehat{\beta}$ via (7), where the tuning parameter ${k}$ is selected via 10-fold cross-validation.
8:if $\widehat{q}+\widehat{k}<p$ then output $\widehat{\beta}$;
9:else $\beta$ is not identifiable.
4.2 Theoretical properties
In this section, we study the theoretical properties of the estimator $\widehat{\beta}$ in Algorithm 2. We consider two paradigms: (1) low-dimensional settings where the dimension of exposure, $p$, is fixed; and (2) high-dimensional settings where $p$ grows with the sample size $n$. For the former, we show that under mild regularity conditions, $\widehat{\beta}$ is $\sqrt{n}$-consistent. For the latter, we show that under mild regularity conditions, $\widehat{\beta}$ achieves a lasso-type error bound. We also demonstrate variable selection consistency in both scenarios. In our theoretical results, we do not require $\widehat{\Lambda}=\Lambda$. We only need certain norms of $\widehat{\Lambda}$ to be consistent with $\Lambda$, which can be achieved by classical estimators.
We first introduce assumptions for the low-dimensional case.
Assumption 1.
(Assumptions for fixed $p$)
  1. B1
    All coefficients ${\Lambda}$, $\beta$, and the function $g(\cdot)$ in models (1) and (3) are fixed and do not change as $n$ goes to infinity.
  2. B2
    $U_{i}$, $\epsilon_{x,i}$, and $\epsilon_{y,i}$ are independent random draws from the joint distribution of $(U,\epsilon_{x},\epsilon_{y})$ such that $E(\epsilon_{x})=\bm{0}$, $E(U)=\bm{0}$, $\text{Cov}(\epsilon_{x})=D$, $\text{Cov}(U)=I_{q}$, and $(U,\epsilon_{x},\epsilon_{y})$ are mutually independent. Furthermore, assume that $\text{Var}({\epsilon}_{y})={\sigma}^{2}$ and $\max_{1\leq j\leq p}\text{Var}(X_{j})=\sigma_{x}^{2}$; these parameters are fixed and do not change as $n$ goes to infinity.
  3. B3
    For the maximum likelihood estimator $\widehat{\Lambda}$, there exists an orthogonal matrix ${O}\in\mathbb{R}^{q\times q}$ such that $\|\widehat{\Lambda}-\Lambda O\|_{2}=O_{p}(1/\sqrt{n})$.
  4. B4
    Let ${\Sigma}_{\widetilde{X}}=\text{Cov}(\widetilde{X})$. We assume $\min_{\theta\in\mathbb{R}^{p},0<\|\theta\|_{0}\leq 2s}\frac{\theta^{\mathrm{\scriptscriptstyle T}}\Sigma_{\widetilde{X}}\theta}{\|\theta\|_{2}^{2}}>c$ for some positive constant $c$.
Conditions B1B2 are standard assumptions for the low-dimensional setting. Given Condition A2, Condition B3 assumes that the estimator for factor loading is root-$n$ consistent. Condition B4 is the population version of the sparse eigenvalue condition (Raskutti et al., 2011, Assumption 3(b)).
Under these conditions, $\widehat{\beta}$ is root-$n$ consistent and achieves consistency in variable selection.
Theorem 2.
Under Conditions A1A4 and B1B4, if the tuning parameter satisfies $\widehat{k}=s$, the following holds:
  1. 1.
    ($\ell_{1}$-error rate) $\|\widehat{\beta}-\dot{\beta}\|_{1}=O_{p}(n^{-1/2}).$
  2. 2.
    (Variable selection consistency) Let $\mathcal{A}=\{j:\dot{\beta}_{j}=0\}$ and $\widehat{\mathcal{A}}=\{j:\widehat{\beta}_{j}=0\}$. Then $\mathbb{P}(\widehat{\mathcal{A}}={\mathcal{A}})\rightarrow 1$ as $n\rightarrow\infty.$
In Theorem 2, it is assumed that $\widehat{k}=s$. This is a standard condition in the $\ell_{0}$-optimization literature (e.g. Raskutti et al., 2011; Shen et al., 2013).
Next, we consider the high-dimensional case and demonstrate that our estimator exhibits properties similar to those of standard regularized estimators in the high-dimensional statistics literature, including a lasso-type error bound and consistency in variable selection. We impose the following regularity conditions.
Assumption 2.
(Assumptions for diverging $p$)
  1. C1
    $sq^{2}\log(p)\log(n)/n\rightarrow 0$, $n=O(p)$, and $q+\log(p)\lesssim\sqrt{n}$, where $x\lesssim y$ means there exist a constant $C$ such that $x\leq Cy$.
  2. C2
    The expectation $\gamma:=\mathbb{E}(Ug(U))\in\mathbb{R}^{q}$, the variance $\sigma_{g}^{2}=\text{Var}(g(U))$, and the covariance $\Gamma:=\text{Var}(Ug(U))\in\mathbb{R}^{q\times q}$ exist. For a matrix $M$, let $\lambda_{\max}(M)$ and $\lambda_{\min}(M)$ denote the maximum and minimum eigenvalues of $M$. There exist positive constants $C_{1}$, $C_{2}$, and $C_{3}$ such that $0<C_{1}\leq\min\{\lambda_{\min}(D),\lambda_{\min}(\Lambda^{\mathrm{\scriptscriptstyle T}}\Lambda/p)\}\leq\max\{\lambda_{\max}(D),\lambda_{\max}(\Lambda^{\mathrm{\scriptscriptstyle T}}\Lambda/p)\}\leq C_{2}<\infty$, and $\max\{\|\gamma\|_{2},\;\sigma_{g}^{2},\;\text{Trace}(\Gamma)\}\leq C_{3}$.
  3. C3
    Assume the random variables in models (1) and (3) satisfy $E(X)=\bm{0}$, $E(U)=\bm{0}$, and $\text{Cov}(U)=I_{q}$. We also assume $\epsilon_{y}$ is independent of $(X,U)$ and ${\epsilon_{x}}$ is independent of $U$. Furthermore, assume ${\epsilon}_{y}$, $\epsilon_{x,j}$, and $X_{j}$ are sub-Gaussian random variables with sub-Gaussian parameters ${\sigma}^{2}$, $\widetilde{\sigma}_{j}^{2}$, and $\sigma_{j}^{2}$. The parameters satisfy ${\sigma}^{2}\leq C_{4}$, and $C_{5}\leq\widetilde{\sigma}_{j}^{2},\sigma_{j}^{2}\leq C_{6}$ for some constants $C_{4},C_{5},C_{6}>0$.
  4. C4
    There exist positive constants $C_{7}$ and $C_{8}$ such that $\underset{i\in\mathcal{A}}{\min}|\dot{\beta}_{i}|\geq n^{C_{7}-1/2}$ and $s^{2}(q+1)^{2}\log{p}\leq n^{2C_{7}-C_{8}}$.
Condition C1 allows the number of exposures $p$ to grow exponentially with the sample size, while the number of latent confounders $q$ grows at a slower polynomial rate. Condition C2 is a standard assumption in high-dimensional factor analysis (Fan et al., 2013; Shen et al., 2016) for loading identification. Condition C3 assumes that the exposures $X_{j}$ are sub-Gaussian, and the noise level is bounded. Condition C4 is a standard assumption on minimum signal strength.
Theorem 3.
Assuming that Conditions A1A4 and C1C3 hold, and the tuning parameter $\widehat{k}=s$, we have:
  1. 1.
    ($\ell_{1}$-error rate) $\|\widehat{\beta}-\dot{\beta}\|_{1}=O_{p}\left(s(q+1)\sqrt{\frac{\log(p)}{n}}\right).$
  2. 2.
    (Variable selection consistency) Under Condition C4, $\mathbb{P}(\widehat{\mathcal{A}}={\mathcal{A}})\rightarrow 1$ as $n\rightarrow\infty.$
Remark 3.
The first part of Theorem 3 differs from Theorem 1 in Ćevid et al. (2020). Their theoretical result relies on the following linear model and decomposition:
$Y=X^{\mathrm{\scriptscriptstyle T}}\beta+U^{\mathrm{\scriptscriptstyle T}}\gamma+\epsilon_{y}=X^{\mathrm{\scriptscriptstyle T}}\beta+X^{\mathrm{\scriptscriptstyle T}}b+(U^{\mathrm{\scriptscriptstyle T}}\gamma-X^{\mathrm{\scriptscriptstyle T}}b)+\epsilon_{y},$
where $b$ is the best linear predictor of $U$ from $X$. Their result depends on (i) $\|b\|_{2}=O_{p}\left(\frac{1}{\sqrt{p}}\right)$ and (ii) the term $(U^{\mathrm{\scriptscriptstyle T}}\gamma-X^{\mathrm{\scriptscriptstyle T}}b)+\epsilon_{y}$ being independent of $X$ under their joint Gaussian assumption. However, these assumptions may not hold in our model (3), where $g(U)$ is an unknown function.
4.3 Extension to nonlinear settings
In this section, we extend the SIV method to address scenarios where the treatment $X$ has nonlinear effects on the outcome $Y$. Revisiting model (3), we remove the assumption of linearity, allowing both the treatment $X$ and the unmeasured confounder $U$ to influence the outcome $Y$ through nonlinear relationships:
$Y=f(X;\beta)+g(U)+\epsilon_{y}.$
In this model, the treatment influences the outcome through the nonlinear causal function $f(\cdot;\beta)$, with $\beta\in\mathbb{R}^{p}$ as the parameter of interest. Our focus is on estimating the parameter $\beta$.
The key observation is that, under model (1), the synthetic instruments $\text{SIV}\in\mathbb{R}^{p-q}=B_{\Lambda^{\perp}}^{\mathrm{\scriptscriptstyle T}}\epsilon_{X}$ are linear combinations of $\epsilon_{X}$, which are independent of both $g(U)$, for any measurable $g$, and $\epsilon_{Y}$. Consequently, we have the following vector equation:
$\mathbb{E}\{SIV(Y-f(X;\beta))\}=\mathbb{E}\{SIV(g(U)+\epsilon_{y})\}=0,$
assuming $\beta$ is set to its true value. Following this, we define the population GMM loss:
$G(\beta)=\left\|\mathbb{E}\{SIV(Y-f(X;\beta))\}\right\|_{2}^{2}.$ (8)
For the function $f(X;\beta)$, let $\partial f(X;\beta)/\partial\beta$ denote the $p\times 1$ column vector $\left({\partial f(X;\beta)}/{\partial\beta_{1}},\ldots,{\partial f(X;\beta)}/{\partial\beta_{p}}\right)^{\mathrm{\scriptscriptstyle T}}$ and let $\partial f(X;\beta)/\partial\beta^{\mathrm{\scriptscriptstyle T}}$ denote the $1\times p$ row vector $\left({\partial f(X;\beta)}/{\partial\beta_{1}},\ldots,{\partial f(X;\beta)}/{\partial\beta_{p}}\right)$. We make the following assumptions, denoted as D1 and D2, which are generalizations of conditions $A1$ and $A4$ in the nonlinear setting.
  1. D1
    (Invertibility) The matrix $\mathbb{E}(X{\partial{f(X;\beta)}}/{\partial{\beta^{\mathrm{\scriptscriptstyle T}}}})\in\mathbb{R}^{p\times p}$ is invertible, and any $q\times q$ submatrix of $\mathbb{E}^{-1}(X{\partial{f(X;\beta)}}/{\partial{\beta^{\mathrm{\scriptscriptstyle T}}}})\Lambda$ is also invertible for any $\beta$.
  2. D2
    (Nonlinear plurality rule) Let $C^{*}$ be a subset of $\{1,2,\ldots,p\}$ with cardinality $q$ and $\dot{\beta}_{C^{*}}\neq 0$, and let the synthetic GMM obtained by assuming $\beta_{{C^{*}}}=0$ be $\widetilde{\beta}^{C^{*}}=\underset{\beta\in\mathbb{R}^{p}:\beta_{C^{*}}=0}{\operatorname*{arg\,min}}G(\beta).$ The plurality rule assumes that $\widetilde{\beta}^{{C^{*}}}$ is uniquely defined and that $\max\limits_{\beta\in\mathbb{R}^{p}}|\{C^{*}:\widetilde{\beta}^{C^{*}}=\beta\}|\leq q.$
The following theorem states that $\beta$ is identifiable using synthetic GMM in a manner parallel to Theorem 1, but within a nonlinear setting.
Theorem 4 (Synthetic Generalized Method of Moments).
Suppose that models (1), (3) and conditions A2, D1, and D2 hold.
  1. 1.
    If A3 holds, then $\dot{\beta}$ is identifiable via $\dot{\beta}=\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;G(\beta)$, subject to $||\beta||_{0}<p-q$.
  2. 2.
    If A3 fails, then $\dot{\beta}$ is not identifiable, and for all $\widetilde{\beta}\in\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;G(\beta)$, we have $||\widetilde{\beta}||_{0}\geq p-q$.
We discuss how to estimate $\beta$ in the finite sample setting. We replace the expectation in equation (8) with the empirical average. Specifically, we consider the following loss function:
$G_{n}(\beta)=\left(\frac{1}{n}\sum^{n}_{i=1}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right)^{\mathrm{\scriptscriptstyle T}}W\left(\frac{1}{n}\sum^{n}_{i=1}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right),$
where $W\in\mathbb{R}^{(p-q)\times(p-q)}$ is a weight matrix. The oracle weight matrix that achieves the highest efficiency in GMM is the inverse of $\mathbb{E}\left[\text{SIV}(Y-f(X;\dot{\beta}))^{2}\text{SIV}^{\mathrm{\scriptscriptstyle T}}\right]=\widetilde{\sigma}^{2}\text{Cov}(\text{SIV})$, where $\widetilde{\sigma}^{2}=\text{Var}(g(U)+\epsilon_{y})$ (Burgess et al., 2017). In practice, we estimate the weight matrix $W$ using the inverse of empirical covariance: $\widehat{\text{Cov}}^{-1}(\widehat{SIV})$.
Finally, let $\bm{X}\in\mathbb{R}^{n\times p}$, $\bm{Y}\in\mathbb{R}^{n\times 1}$, $f(\bm{X};\beta)\in\mathbb{R}^{n\times 1}$, and $\bm{SIV}\in\mathbb{R}^{n\times(p-q)}$ denote the relevant random matrices in the finite sample setting. We consider the following optimization problem:
$\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\|\bm{SIV}({{\bm{SIV}^{\mathrm{\scriptscriptstyle T}}\bm{SIV}}})^{-1}{{\bm{SIV}^{\mathrm{\scriptscriptstyle T}}(\bm{Y}-f(\bm{X};\beta))}}\right\|^{2}_{2}\quad\text{subject to }\|\beta\|_{0}\leq k.$ (9)
The unconstrained optimization in (9) is also referred to as the nonlinear two-stage least squares estimator in the literature (Amemiya, 1974). Under the linear setting where $f(\bm{X};\beta)=\bm{X}\beta$, (9) has a similar form to (7). Optimization (9) can be accelerated using a splicing approach, which is designed to efficiently solve the best subset selection problem (Zhu et al., 2020; Zhang et al., 2023).
5 Simulation studies
In this section, we evaluate the numerical performance of the proposed SIV method and compare it with other methods across various scenarios. First, we assess the performance of the SIV method under a linear outcome model in Section 5.1. Next, we evaluate the algorithm in the context of a nonlinear outcome model in Section 5.2. In Section LABEL:sec:MS of the supplementary material, we explore an alternative moment selection estimator inspired by the work of Andrews (1999) and compare it with our proposed estimator.
5.1 Simulation studies with a linear outcome model
We begin by evaluating the SIV estimator within a linear outcome model. The model is defined by $f(X;\beta)=X^{\mathrm{\scriptscriptstyle T}}\beta$ and $g(U)=U^{\mathrm{\scriptscriptstyle T}}\gamma$. In our simulations, we let $q=3$, $s=5$, and $\beta=(1,1,1,1,1,0,\ldots,0)^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{p}$. Each element in $\Lambda_{j,k}$ and $\gamma_{k}$ is independently generated from $\text{Uniform}(-1,1)$ for $j=1,\ldots,p$, and $k=1,\ldots,q$. The hidden variables $U_{i,k}$ follow i.i.d. standard normal distributions for $i=1,\ldots,n$ and $k=1,\ldots,q$. The random errors are generated as $\epsilon_{x}\sim\mathbb{N}(0,\sigma^{2}_{x}I_{p})$ and $\epsilon_{y}\sim\mathbb{N}(0,\sigma^{2})$, where $\sigma_{x}=2$ and $\sigma=5$. We evaluate the performance of our method under the following two settings: (i) Low-dimensional cases: $p=100$ and $n\in\{200,600,1000,\ldots,5000\}$; (ii) High-dimensional cases: $n=500$ and $p\in\{500,750,1000,\ldots,3000\}$. All simulation results are based on 1000 Monte Carlo runs. The data-generating mechanism is designed to mimic certain features of the real application in Section 6; see Section LABEL:sec:comparison_sim&data in the supplementary material for more details.
We compare the following methods in our simulations.
  1. 1.
    (SIV) We implement Algorithm 2 and determine $\widehat{q}$ using the method proposed by Onatski (2010). A detailed discussion of Onatski (2010)’s method is provided in Section LABEL:rem:Onatski of the supplementary material. For cases where $p\leq 30$, we employ a full best subset selection routine to solve the $\ell_{0}$-optimization problem. When $p>30$, we utilize the adaptive best subset selection method implemented in the abess function in R.
  2. 2.
    (Lasso, Tibshirani, 1996): We implement the lasso using the glmnet function in R, with the tuning parameter selected via 10-fold cross-validation.
  3. 3.
    (Trim, Ćevid et al., 2020): We implement Ćevid et al. (2020)’s method using the code available from https://github.com/zijguo/Doubly-Debiased-Lasso.
  4. 4.
    (Null, Miao et al., 2023): For the low-dimensional settings, we implement Miao et al. (2023)’s method using the code available from https://www.tandfonline.com/doi/suppl/10.1080/01621459.2021.2023551. For the high-dimensional settings, their method cannot be applied directly because $\xi$ in their method cannot be solved by ordinary least squares. Therefore, we replace ordinary least squares with the lasso, where the tuning parameter is selected by 10-fold cross-validation.
  5. 5.
    (IV-Lasso): Motivated by a reviewer’s comment, we consider the following “IV-Lasso” estimator. In the first step, we use Lasso to select active predictors and define $\widehat{\mathcal{A}}=\{j:\widetilde{\beta}_{j}\neq 0\}$, where $\widetilde{\beta}$ is the Lasso estimator described above. In the next step, we use $\widehat{SIV}$ as instrumental variables to estimate the treatment effects of $X_{\widehat{\mathcal{A}}}$ on $Y$ using the ivreg package, denoted as $\widehat{\beta}^{\;\text{IV-Lasso}}_{\widehat{\mathcal{A}}}$, and set $\widehat{\beta}^{\;\text{IV-Lasso}}_{\widehat{\mathcal{A}}^{c}}=0$.
(a) $p=100$, and $n$ varies from $200$ to $5000$.
(b) $n=500$, and $p$ varies from $500$ to $3000$.
[…]
(a) $p=100$, and $n$ varies from $200$ to $5000$.
(b) $n=500$, and $p$ varies from $500$ to $3000$.
[…]
We present the $\ell_{1}$-estimation errors of the five methods in Figures 2. In low-dimensional cases, as illustrated in Figure 2(a), the bias of the Lasso estimator stabilizes as the sample size grows. The Lasso estimator does not account for unmeasured confounding and is, therefore, not expected to perform well. The Lasso, Trim, and Null methods exhibit considerable $\ell_{1}$-estimation bias relative to other methods. Compared to IV-Lasso, our SIV estimator demonstrates superior performance. This advantage arises from the $\ell_{0}$-optimization employed in the SIV method, which enables more accurate variable selection, as illustrated in Figure 3(a) later.
For the high-dimensional settings, we can see from Figure 2(b) that our SIV method consistently outperforms the comparison methods. As discussed in Section 1, the true correlations between $X$ and $Y$ are non-sparse. This explains the large estimation errors of the Lasso method. The Null method exhibits even larger biases than the Lasso method. The Trim method, designed specifically for high-dimensional settings, outperforms both the Lasso and Null methods. However, our estimator still shows a much smaller bias compared to the Trim estimator.
We also observe from Figure 2(b) that the IV-Lasso estimator underperforms compared to the naive Lasso estimator in high-dimensional settings. This underperformance occurs because the naive Lasso method applies $\ell_{1}$ penalization, which drives the coefficients of incorrectly selected $\widehat{\beta}_{i}$ towards zero. In contrast, the second step of IV-Lasso negates this shrinkage, causing the coefficients of incorrectly selected $\widehat{\beta}_{i}$ to diverge from zero, thereby increasing estimation bias.
[…]
We further evaluate the performance of our proposed algorithm in settings with non-diagonal $\text{Cov}(\epsilon_{x})$. We generate $\epsilon_{x,i}\sim\mathbb{N}(0,D)$. For the low-dimensional setting, we randomly select 20 pairs of $i,j\in\{1,2,\ldots,p\}$ and set $D_{i,j}=D_{j,i}=1$. The list of pairs is provided in Section LABEL:sec:detail:nondiagonal of supplementary material. In the high-dimensional settings, we set $D_{i,j}=4\times 0.3^{|i-j|}$. All other aspects of the data-generating mechanism remain unchanged. In the low-dimensional scenario, we use the Robust Principal Component Analysis method (Candès et al., 2011) to estimate the low-dimensional structure of the covariance matrix and factor loadings. For the high-dimensional scenario, we directly use the Principal Component Analysis method to estimate factor loadings. These are implemented using the R packages rpca and prcomp, respectively.
The simulation results, presented in Figures 4 and 5
illustrate the $\ell_{1}$-estimation errors and false discovery rates for various methods. These methods exhibit similar trends in $\ell_{1}$-error performance as observed in the previous setting with uncorrelated errors. The Trim method shows a slight improvement, achieving lower $\ell_{1}$-errors and false discovery rates compared to the previous simulation results. The performance of the SIV method remains consistent with its performance when $D$ is a diagonal matrix. Additionally, our method continues to outperform the other comparison methods.
(a) $p=100$, and $n$ varies from $200$ to $5000$.
(b) $n=500$, and $p$ varies from $500$ to $3000$.
[…]
(a) $p=100$, and $n$ varies from $200$ to $5000$.
(b) $n=500$, and $p$ varies from $500$ to $3000$.
[…]
5.2 Simulation studies with non-linear outcome models
We then evaluate the performance of our proposed estimator (9) with nonlinear outcome models. We set $q=2$, $s=2$, and $p=10$. We consider two different settings for $f(X)$. In the first setting, $f(X;\beta)=\sum^{10}_{j=1}X_{j}^{3}\beta_{j}$ with $\beta=(0.3,0.3,0,0,\ldots,0)^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{10}$ and $g(U)=U_{1}^{3}\gamma_{1}+U_{2}^{3}\gamma_{2}$. In the second setting, $f(X;\beta)=\exp(X^{\mathrm{\scriptscriptstyle T}}\beta)$ and $g(U)=(U^{3})^{{\mathrm{\scriptscriptstyle T}}}\gamma$. Each element in $\Lambda_{j,k}$ and $\gamma_{k}$ is independently generated from $\mathbb{N}(0,1)$ for $j=1,\ldots,p$ and $k=1,\ldots,q$. The hidden variables $U_{i,k}$ follow i.i.d. standard normal distributions for $i=1,\ldots,n$ and $k=1,\ldots,q$. The random errors are generated by $\epsilon_{x}\sim\mathbb{N}(0,\sigma^{2}_{x}I_{p})$ and $\epsilon_{y}\sim\mathbb{N}(0,\sigma^{2})$, where $\sigma_{x}=2$ and $\sigma=1$. We evaluate the performance of our estimators with $n\in\{1000,\ldots,5000\}$. All simulation results are based on 1000 Monte Carlo runs.
For comparison, we did not implement the other methods in Section 5.1, as they are not designed for non-linear outcome models. Instead, we consider a popular method for addressing endogeneity in high-dimensional settings (e.g. Wang and Blei, 2019; Ouyang et al., 2023; Fan et al., 2024), which first estimates the unmeasured confounders $U$ and then directly adjusts for its estimate $\widehat{U}$ in the regression modeling. Specifically, we compare the proposed method with the following two variations of the so-called U-hat method, with the tuning parameter $k$ selected using 10-fold cross-validation in all cases:
  1. 1.
    (SIV): We obtain $\widehat{\beta}$ from (9). The $\ell_{0}$-optimization is accelerated using the splicing technique (Zhang et al., 2023).
  2. 2.
    (U-hat1): First, we obtain $\widehat{\bm{U}}\in\mathbb{R}^{n\times q}$ using the equation $\widehat{\bm{U}}=\bm{X}\widehat{\text{Cov}}(X)^{-1}\widehat{\Lambda}$. Next, we obtain $\widehat{\beta}$ by solving the following optimization problem: $\widehat{\beta}=\underset{\beta\in\mathbb{R}^{p},\gamma\in\mathbb{R}^{q}}{\operatorname*{arg\,min}}\|\bm{Y}-f(\bm{X};\beta)-\widehat{\bm{U}}\gamma\|_{2}^{2}\quad\text{subject to}\quad\|\beta\|_{0}\leq k.$
  3. 3.
    (U-hat2): We first obtain $\widehat{\bm{U}}\in\mathbb{R}^{n\times q}$ similarly. We then obtain $\widehat{\beta}$ by solving the following optimization problem: $\widehat{\beta}=\underset{\beta\in\mathbb{R}^{p},\gamma\in\mathbb{R}^{q}}{\operatorname*{arg\,min}}\|\bm{Y}-f(\bm{X};\beta)-\widehat{\bm{U}}^{3}\gamma\|_{2}^{2}\quad\text{subject to}\quad\|\beta\|_{0}\leq k,$ where $\widehat{\bm{U}}^{3}\in\mathbb{R}^{n\times q}$ is defined as $\{\widehat{\bm{U}}^{3}\}_{i,j}=\{\widehat{\bm{U}}_{i,j}\}^{3}$. Note this method assumes knowledge of the specific form of $g(U)$, which is generally not available in practice.
Figure 6 graphically presents the $\ell_{1}$-estimation error for all estimators. The results show that U-hat1 and U-hat2 have similar performance, with bias stabilizing as the sample size increases. In contrast, the bias of our estimator decreases with larger samples, indicating that our method consistently estimates the causal parameter $\beta$ in nonlinear outcome models.
Remark 4.
Under the linear models in Section 5.1, results from the U-hat (including U-hat1 and U-hat2) and SIV methods coincide numerically. While the $U$-hat method has been proposed in the literature (e.g., Wang and Blei, 2019), its validity has been widely challenged (e.g., Ogburn et al., 2019; Grimmer et al., 2023). From this perspective, our results may be viewed as a justification for the $U$-hat method in the special and unrealistic case of linear outcome model (2). In general, however, our method is fundamentally different from the U-hat method.
The U-hat method relies on regression adjustment, which assumes a specified outcome-confounder relationship. If confounders $U$ were observed, then, at least in principle, this relationship could be estimated from the data. But if $U$ is unobserved, then specifying the functional form of $g$ depends entirely on subject-matter knowledge, which is rarely available. In Section LABEL:sec:Uhat of the supplementary material, we explain in detail why the $U$-hat method fails under the general model $Y=f(X;\beta)+g(U)+\epsilon_{y}$.
In contrast, our approach constructs instrumental variables, which are agnostic to the form of unmeasured confounding even if the causal relationship is non-linear. Because of this, we believe it is a more robust approach in this setting. To our knowledge, we are the first to apply the IV framework to this problem, offering a method that accommodates nonlinear causal effects and arbitrary dependence on unmeasured confounders $U$. See Section LABEL:sec:Uhat in the supplementary materials for more discussions.
(a) Non-linear setting 1.
(b) Non linear setting 2.
[…]
6 Real data application
To further illustrate the proposed synthetic instrumental variable method, we reanalyzed a mouse obesity dataset described by Wang et al. (2006). The study involved a cross of 334 mice between the C3H strain and the susceptible C57BL/6J (B6) strain on an ApoE-null background, which were fed a Western diet for 16 weeks. The dataset includes genotype data on 1,327 SNPs, gene expression profiles of 23,388 liver tissue genes, and clinical information such as body weights. Lin et al. (2015) previously analyzed this dataset using regularized methods for high-dimensional instrumental variable regression, treating the SNPs as potential instruments, the gene expressions as treatments, and identifying 17 genes likely to affect mouse body weight. Gleason et al. (2021) also discussed controversies surrounding the use of SNPs as instruments for estimating the effects of gene expression. Miao et al. (2023) applied their method to estimate the causal effects associated with these 17 genes. However, their approach cannot be used to estimate effects associated with the full set of 23,388 genes, as it only accommodates low-dimensional exposures. In our analysis, we use the same dataset to identify the genes that influence mouse body weight and to estimate the magnitude of these effects. Notably, our method does not depend on genotype data or other instrumental variables and can handle high-dimensional exposures.
We followed the procedure described in Lin et al. (2015) to preprocess the dataset. Genes with a missing rate greater than 0.1 were removed, and the remaining missing gene expression data were imputed using nearest neighbor averaging (Troyanskaya et al., 2001). We also removed genes that could not be mapped to the Mouse Genome Database (MGD) and those with a standard deviation of gene expression levels less than $0.1$. A marginal linear regression model was fitted between the mice’s body weight and sex to subtract the estimated sex effect from the body weight, adjusting for the effect of sex. The fitted residual was used as the outcome $Y$, and the gene expression levels were centered and standardized as multiple treatments. After data cleaning and merging gene expression and clinical data, the final dataset comprised $p=2819$ genes from $n=306$ mice ($154$ female and $152$ male).
The estimator proposed by Onatski (2010) suggests that there are three unobserved latent factors, so we applied our method with $\widehat{q}=3$, using a linear outcome model as the working model. Five genes were found to affect mouse body weight. Under the plurality rule A4, we conclude that the causal effects are identifiable as $\widehat{q}+\widehat{s}=8\ll p=2819$. Specifically, our analysis suggests that increasing the gene expression levels of $Igfbp2$, $Rab27a$, $Dct$, $Ankhd1$, and $Gck$ by one standard deviation leads to changes of $-1.98$, $1.88$, $1.43$, $-1.33$, and $1.17$ grams in mouse body weight, respectively.
We compared our approach with six methods: (i) the Lasso method; (ii) the two-stage regularization (2SR) method (Lin et al., 2015), which leverages SNPs as high-dimensional instrumental variables to estimate the causal effect; (iii) the auxiliary variable method (Miao et al., 2023), which focuses on the 17 genes identified by Lin et al. (2015) as having non-zero effects and uses the five SNPs selected by Lin et al. (2015) as auxiliary variables; (iv) the null variable method (Miao et al., 2023), which assumes that more than half of the 17 genes have zero effects on mouse body weight; (v) the Trim method (Ćevid et al., 2020); and (vi) the IV-Lasso method, as detailed in Section 5. Detailed results for these comparison methods are included in Supplementary Material Section LABEL:sec:comparison. The number of active genes found by these methods, defined as genes that have non-zero effects on mouse body weight, is 87, 17, 4, 2, 4, and 14, respectively. Consistent with the simulation results, the Lasso method identifies many more active exposures than our method, and its selected set includes all genes identified by our approach. Compared to the other methods, both the 2SR and auxiliary variable methods rely on additional SNP information.
All methods, except the null variable method, identify the expression of the $Igfbp2$ gene (insulin-like growth factor binding protein 2) as a cause of obesity, which is known to prevent obesity and protect against insulin resistance (Wheatcroft et al., 2007). Additionally, we identify four other genes potentially linked to obesity. The $Rab27a$ gene (Ras-related protein Rab-27A) is involved in insulin granule docking in pancreatic $\beta$ cells, and Rab27a-mutated mice show glucose intolerance after a glucose load (Kasai et al., 2005), suggesting its positive effect on body weight. The $Dct$ gene (Dopachrome Tautomerase) has been associated with obesity and glucose intolerance (Kim et al., 2015), with overexpression observed in the visceral adipose tissue of morbidly obese patients (Randhawa et al., 2009). The $Gck$ gene (Glucokinase) plays a key role in blood glucose recognition, and its overexpression is linked to insulin resistance (Randhawa et al., 2009), which may explain its impact on body weight.
7 Discussion
In this paper, we study how to identify and estimate causal effects with a multi-dimensional treatment in the presence of unmeasured confounding. Our key assumption is a sparse causation assumption, which in many contexts serves as an appealing alternative to the widely adopted sparse association assumption. We develop a synthetic instrument approach to identify and estimate causal effects without the need for collecting additional exogenous information, such as instrumental variables. Our estimation procedure can be formulated as an $\ell_{0}$-optimization problem and can therefore be solved efficiently using off-the-shelf packages.
A distinctive feature of our framework is that it allows the use of Algorithm 2 to consistently test the sparsity condition A3 under the other model assumptions. In practice, however, we observe that this test can be unstable in finite samples, particularly in boundary cases where $p\approx s+q$. Developing more stable tests for the sparsity condition A3 is an interesting avenue for future research.
We have focused on a linear treatment model (1). In a general nonlinear treatment model, where the relationship between treatment $X$ and the unmeasured factor $U$ is nonlinear, nonlinear factor analysis could be used to fit $X=m(U;\alpha)+\epsilon$. However, identifying a function $h(\cdot)$ such that $h(X)$ is independent of $U$ remains a significant challenge. Extending this framework to accommodate nonlinear treatment models is left for future research.
We have also focused on the identification and estimation problems. Assuming that the $\ell_{0}$-penalization procedure (7) accurately selects the true non-zero causal effects, standard M-estimation theory can be used to construct pointwise confidence intervals. However, constructing uniformly valid confidence intervals for the causal parameters remains a challenge, as statistical inference after model selection is typically not uniform (Leeb and Pötscher, 2005). One promising approach is to build on a uniformly valid inference method for the standard $\ell_{0}$-penalization procedure, which, to the best of our knowledge, remains an open problem in the statistical literature.
Acknowledgements
We thank Xin Bing, Zhichao Jiang, Wang Miao, Thomas Richardson, James Robins, Dominik Rothenhäusler, and Xiaochuan Shi for their helpful discussions and constructive comments. We also extend our gratitude to the Editor, Associate Editor, and anonymous referees for their valuable and thoughtful input, which has significantly enhanced the quality of this manuscript.
Supplementary material
The Supplementary Material contains further discussions of the assumptions and additional simulation results. It also includes more examples and proofs of all the theorems and lemmas.

### v3

@@
 Abstract
-In many observational studies, researchers are often interested in studying the effects of multiple exposures on a single outcome.
+In many observational studies, researchers are often interested in the effects of multiple exposures on a single outcome.
 Standard approaches for high-dimensional data, such as the Lasso, assume that the associations between the exposures and the outcome are sparse.
@@
 In this paper, we consider an alternative approach that assumes the causal effects under consideration are sparse.
-We show that with sparse causation, causal effects are identifiable even with unmeasured confounding.
-At the core of our proposal is a novel device called the synthetic instrument, which, in contrast to standard instrumental variables, can be constructed using the observed exposures directly.
-We demonstrate that under the assumption of sparse causation, the problem of causal effect estimation can be formulated as an $\ell_{0}$-penalization problem and can thus be solved efficiently using off-the-shelf software.
-Simulations show that our approach outperforms state-of-the-art methods in both low-dimensional and high-dimensional settings.
+We show that under sparse causation, causal effects are identifiable even with unmeasured confounding.
+Our proposal is built around a novel device called the synthetic instrument, which, in contrast to standard instrumental variables, can be constructed directly from the observed exposures.
+We demonstrate that, under the assumption of sparse causation, the problem of causal effect estimation can be formulated as an $\ell_{0}$-penalization problem and solved efficiently using off-the-shelf software.
+Simulations show that our approach outperforms state-of-the-art methods in both low- and high-dimensional settings.
 We further illustrate our method using a mouse obesity dataset.
@@
 Email: linbo.wang@utoronto.ca
-1Department of Statistics and Data Science, Washington University in St. Louis, St. Louis, Missouri, U.S.A.
+1Department of Mathematics and Statistics, University of Ottawa, Ottawa, Ontario, Canada
 2 Department of Statistical Sciences, University of Toronto, Toronto, Ontario, Canada
@@
 1 Introduction
-Sparsity is a common assumption in modern statistical learning literature, as it allows for variable selection in underlying models and enhances the interpretability of parameter estimates.
-For example, the lasso method (Tibshirani, 1996) assumes that the underlying associations between a single outcome and potentially high-dimensional predictors are sparse.
-In other words, only a small number of predictors have non-zero associations with the outcome.
+Sparsity is a common assumption in the modern statistical learning literature, as it facilitates variable selection in models and enhances the interpretability of parameter estimates.
+For example, the Lasso (Tibshirani, 1996) assumes sparse associations between a single outcome and potentially high-dimensional predictors; in other words, only a small subset of predictors have nonzero associations with the outcome.
 Hastie et al.
-(2009) summarize the philosophy behind these methods as the “bet on sparsity” principle: Use a procedure that performs well in sparse problems, as no procedure performs well in dense problems.
-Methods like the lasso perform well when the associations between the outcome and predictors are sparse and have gained significant popularity in recent decades.
-The “bet on sparsity” principle does not restrict the type of “sparse problems” that can be considered.
-Instead of focusing solely on sparse associations, there is a growing body of literature that emphasizes sparse causation, where only a fraction of the exposures have non-zero causal effects on the outcomes (e.g., Spirtes and Glymour, 1991; Claassen et al., 2013; Wang et al., 2017; Miao et al., 2023; Zhou et al., 2024).
-This assumption is arguably more interpretable and plausible in many data applications.
-For example, suppose one is interested in the relationships between the expression of multiple genes and a phenotype of interest, such as lung cancer.
-Current biological knowledge suggests that not all genes in the human genome may affect lung cancer (e.g., Kanwal et al., 2017).
-However, this does not imply the assumption of sparse association, as unmeasured confounding often leads to spurious correlations between the exposures and the outcome.
-For example, consider a linear structural model (Pearl, 2013) with a $p$-dimensional exposure $X=(X_{1},\ldots,X_{p})^{\mathrm{\scriptscriptstyle T}}$, an outcome $Y$, and a $q$-dimensional latent variable $U$:
+(2009) summarizes the philosophy behind such methods as the “bet on sparsity” principle: use a procedure that performs well in sparse settings, since no procedure performs well in dense ones.
+Methods like the Lasso perform well under sparse associations and, as a result, have gained significant popularity in recent decades.
+Importantly, the “bet on sparsity” principle does not restrict the types of problems to which sparsity may apply.
+Beyond sparse associations, a growing body of literature emphasizes sparse causation, where only a fraction of exposures exert nonzero causal effects on the outcome (e.g., Spirtes and Glymour, 1991; Claassen et al., 2013; Wang et al., 2017; Miao et al., 2023; Zhou et al., 2024).
+This assumption is often more interpretable and plausible in real data applications.
+For example, suppose we are interested in the relationship between gene expression and a phenotype such as lung cancer.
+Biological evidence suggests that only a small proportion of genes may influence the risk of lung cancer (e.g., Kanwal et al., 2017).
+However, this does not imply sparse association, since unmeasured confounding may induce spurious correlations between many genes and the phenotype.
+To illustrate, consider a linear structural model (Pearl, 2013) with a $p$-dimensional exposure vector $X=(X_{1},\ldots,X_{p})^{\mathrm{\scriptscriptstyle T}}$, an outcome $Y$, and a $q$-dimensional latent variable $U$:
 $X$ $=\Lambda U+\epsilon_{x},$ (1)
 $Y$ $=X^{\mathrm{\scriptscriptstyle T}}\beta+U^{\mathrm{\scriptscriptstyle T}}\gamma+\epsilon_{y},$ (2)
-where $\Lambda\in\mathbb{R}^{p\times q}$, $\beta\in\mathbb{R}^{p}$, and $\gamma\in\mathbb{R}^{q}$ are coefficients, and $\epsilon_{x}=(\epsilon_{1},\ldots,\epsilon_{p})^{\mathrm{\scriptscriptstyle T}}$, $\epsilon_{y}$, and $U$ are uncorrelated with each other.
-Under this model, the spurious correlations due to unmeasured confounding, characterized by $\text{Cov}(X)^{-1}\Lambda\gamma$, are typically dense.
-As a result, the association between $X$ and $Y$ is not sparse, even if the causal effect $\beta$ is sparse.
-Identification and estimation of the causal parameter $\beta$ are non-trivial due to the presence of unmeasured confounding by $U$.
-Our contributions in this paper are twofold.
-First, under an additional plurality condition, we establish that the parameter $\beta$ in model (3) is identifiable if and only if $||\beta||_{0}<p-q$.
-Our assumption on the sparsity level is both necessary and sufficient, representing a significant improvement over assumptions previously introduced in the literature; see Section 1.1 for details.
-Remarkably, in contrast to many other identification assumptions in causal inference, it is possible to consistently test the sparsity assumption from data.
-Second, we develop a synthetic two-stage regularized regression approach for estimating $\beta$, with a first-stage ordinary least squares and a second-stage $\ell_{0}$-penalized regression.
-The key technique behind our results is a novel device we call the synthetic instrument.
-Unlike standard instrumental variables, the synthetic instrument is constructed from a subset of exposures, allowing us to identify the causal effects without resorting to exogenous variables.
-Our procedure comes with lasso-type theoretical guarantees in both low- and high-dimensional settings.
+where $\Lambda\in\mathbb{R}^{p\times q}$, $\beta\in\mathbb{R}^{p}$, and $\gamma\in\mathbb{R}^{q}$ are coefficient vectors, and $\epsilon_{x}=(\epsilon_{1},\ldots,\epsilon_{p})^{\mathrm{\scriptscriptstyle T}}$, $\epsilon_{y}$, and $U$ are mutually uncorrelated.
+Under this model, spurious correlations induced by unmeasured confounding, given by $\text{Cov}(X)^{-1}\Lambda\gamma$, are typically dense.
+Consequently, the overall association between $X$ and $Y$ is dense, even when the causal effect $\beta$ itself is sparse.
+Identification and estimation of the causal parameter $\beta$ are nontrivial due to the presence of unmeasured confounding by $U$.
+The contributions of this paper are twofold.
+First, under an additional plurality condition, we establish that the parameter $\beta$ in model (3) is identifiable if and only if $\|\beta\|_{0}<p-q$.
+This sparsity assumption is both necessary and sufficient, representing a significant improvement over conditions previously introduced in the literature; see Section 1.1 for details.
+Remarkably, in contrast to many other identification assumptions in causal inference, this assumption can be consistently tested from data.
+Second, we develop a two-stage synthetic regularized regression approach for estimating $\beta$, with a first stage based on ordinary least squares and a second stage using $\ell_{0}$-penalized regression.
+The key technique behind our results is a novel device, which we term the synthetic instrument.
+Unlike standard instrumental variables, the synthetic instrument is constructed from a subset of exposures, enabling identification of causal effects without requiring exogenous variables.
+Our procedure enjoys Lasso-type theoretical guarantees in both low- and high-dimensional settings.
 1.1 Related works
-Our proposal is related to recent work dealing with multivariate hidden confounding.
+Our proposal is related to recent work on multivariate hidden confounding.
 Ćevid et al.
 (2020) and Guo et al.
-(2022) propose a spectral deconfounding method for estimating $\beta$ under a high-dimensional model.
-Their method assumes a dense confounding model, which is only possible under a high-dimensional regime where $p$ tends to infinity with the sample size and the magnitude of spurious associations tends to zero.
+(2022) propose a spectral deconfounding method for estimating $\beta$ in a high-dimensional model.
+Their method assumes a dense confounding structure, which is feasible only in a high-dimensional regime where $p$ tends to infinity with the sample size and the magnitudes of spurious associations tend to zero.
 Bing et al.
-(2022) consider a more general setup than ours, in which they also allow the outcome to be multivariate.
+(2022) consider a more general setup than the one we study, in which they also allow the outcome to be multivariate.
 However, they aim to identify the projection of $\beta$ onto a related space rather than the causal parameter $\beta$ itself.
@@
 Under this assumption, they not only identify the effect of $X$ on $Y$ but also recover the covariance among components of $X$ conditional on $U$.
-The estimation problem of $\beta$ can be framed within the context of causal inference with unmeasured confounding.
-Currently, the most popular approach in practice is the instrumental variable (IV) framework, which uses information from an exogenous variable known as an IV to identify causal effects.
-(e.g., Angrist et al., 1996; Wang and Tchetgen Tchetgen, 2018).
-Another approach that has gained attention recently is the proximal causal inference framework (Tchetgen Tchetgen et al., 2024), which uses information from ancillary variables known as negative control exposures and outcomes to remove bias due to unmeasured confounding.
+The estimation problem for $\beta$ can be framed within the context of causal inference with unmeasured confounding.
+Currently, the most popular approach in practice is the instrumental variable (IV) framework, which uses information from an exogenous variable known as an IV to identify causal effects (e.g., Angrist et al., 1996; Wang and Tchetgen Tchetgen, 2018; Pfister and Peters, 2022).
+Another approach that has gained attention recently is the proximal causal inference framework (Tchetgen Tchetgen et al., 2024), which uses information from ancillary variables, known as negative control exposures and outcomes, to remove bias due to unmeasured confounding.
 Compared with these frameworks, our approach does not rely on the collection of additional ancillary variables, which can be challenging in many practical settings.
 Instead, we rely on the availability of multiple exposures and the sparsity assumption for identification and estimation.
-Recently, there has been a strand of literature that seeks to identify the causal effects of multiple exposures.
-Wang and Blei (2019) popularized this setting by proposing the so-called deconfounder method, which first obtains an estimate $\widehat{U}$ for the unmeasured confounder and then adjusts for $\widehat{U}$ using standard regression methods.
-However, it has been pointed out that under this setting, without further assumptions, the causal effect $\beta$ is not identifiable (D’Amour, 2019; Ogburn et al., 2020).
+Recently, a strand of literature has sought to identify the causal effects of multiple exposures.
+Wang and Blei (2019) popularized this setting by proposing the so-called deconfounder method, which first obtains an estimate $\widehat{U}$ of the unmeasured confounder and then adjusts for $\widehat{U}$ using standard regression methods.
+However, it has been pointed out that in this setting, without further assumptions, the causal effect $\beta$ is not identifiable (D’Amour, 2019; Ogburn et al., 2020).
 Kong et al.
@@
 (2023) consider a similar setting to (1) and (3), showing that the causal effect is identifiable if $\|\beta\|_{0}\leq(p-q)/2$.
-Their sparsity constraint is significantly stronger than ours, especially in cases where the number of exposures is large relative to the number of latent confounders.
+Their sparsity constraint is significantly stronger than ours, especially when the number of exposures is large relative to the number of latent confounders.
 Miao et al.
-(2023) also develop a robust linear regression-based estimator for estimating $\beta$.
-In contrast to our estimator, their estimator is only consistent in the low-dimensional regime where $p$ is fixed and $\|\beta\|_{0}\leq p/2-q+1$.
-Furthermore, their estimator for $\beta$ is not sparse and, therefore, cannot be used for selecting treatments with non-zero effects.
+(2023) also develop a robust linear regression-based estimator for $\beta$.
+In contrast to our estimator, their estimator is consistent only in the low-dimensional regime where $p$ is fixed and $\|\beta\|_{0}\leq p/2-q+1$.
+Furthermore, their estimator for $\beta$ is not sparse and therefore cannot be used for selecting treatments with nonzero effects.
 Our results also connect to recent literature on multiply robust causal identification (e.g., Sun et al., 2023), as we show identification in the union of many causal models.
-This contrasts with the rich literature on multiply robust estimators under the same causal model (e.g., Wang and Tchetgen Tchetgen, 2018) and improved doubly robust estimators that are consistent under multiple working models for two components of the likelihood (e.g., Han and Wang, 2013).
+This contrasts with the extensive literature on multiply robust estimators under the same causal model (e.g., Wang and Tchetgen Tchetgen, 2018) and on improved doubly robust estimators that are consistent under multiple working models for two components of the likelihood (e.g., Han and Wang, 2013).
 1.2 Outline of this paper
@@
 In Section 2, we introduce the setup and background.
-In Section 3, we introduce our identification strategy using the synthetic instrument method.
+In Section 3, we describe our identification strategy using the synthetic instrument method.
 In Section 4, we present our estimation procedure and provide theoretical justifications.
-We also discuss extensions to non-linear outcome models.
-Simulation studies in Section 5 compare our proposal with several state-of-the-art methods in terms of their finite-sample performance.
+We also discuss extensions to nonlinear outcome models.
+Simulation studies in Section 5 compare our proposal with several state-of-the-art methods in finite-sample performance.
 In Section 6, we apply our method to mouse obesity data.
 We conclude with a brief discussion in Section 7.
-The proposed method is implemented in an R package available at https://github.com/dingketang/syntheticIV.
+The proposed method is implemented in an R package, available at https://github.com/dingketang/syntheticIV.
 2 Framework, notation, and identifiability
@@
 In Section 4.3, we will consider a nonlinear treatment effect model, where the relationship between treatment $X$ and outcome $Y$ is represented by a potentially nonlinear function $f(X;\beta)$.
-We consider both low- and high-dimensional settings where $p$ may be smaller or larger than the sample size $n$.
-Assume $\dot{\beta}$ is the true value of $\beta$ in model (3).
+We consider both low- and high-dimensional settings, where $p$ may be smaller or larger than the sample size $n$.
+Let $\dot{\beta}$ denote the true value of $\beta$ in model (3).
 Without loss of generality, we assume all the variables in (1) and (3) are centered, $\text{Cov}(U)=I_{q}$, and $\mathbb{E}(g(U))=0$.
@@
 However, it may be relaxed in our setting.
-For example, if some treatment effects are not confounded after normalization, such that certain rows of $\text{Cov}^{-1}(X)\Lambda$ are zero, then although Condition A1 is violated, the proposed method can still be used to identify and estimate the treatment effects.
-See Section LABEL:sec:discussA1 of the supplementary material for more details.
-Condition A2 has been discussed extensively in the factor model literature. One classical result is Proposition 1 below, which is a direct corollary of Anderson and Rubin (1956, Theorem 5.1).
+For example, if certain treatment effects are unconfounded after normalization, so that specific rows of $\text{Cov}^{-1}(X)\Lambda$ are zero, then even though Condition A1 is violated, our proposed method can still be used to identify and estimate the treatment effects.
+See Sections LABEL:sec:discussA1 and LABEL:sec:weakivandidentification of the supplementary material for further details on how the algorithm identifies treatment effects under relaxed versions of A1.
+Condition A2 has been discussed extensively in the factor model literature. One classical result is Proposition 1, which is a direct corollary of Anderson and Rubin (1956, Theorem 5.1).
 Proposition 1.
 Under models (1), (3), and Condition A1, if $p\geq 2q+1$ and $D=\text{Cov}(\epsilon_{x})$ is a diagonal matrix, then $\Lambda$ is identifiable up to a rotation.
-We note that the condition that $D$ is a diagonal matrix is a classical assumption in the factor analysis literature.
+We note that the condition that $D$ be a diagonal matrix is a classical assumption in the factor analysis literature.
 However, Condition A2, and hence our algorithm, may still hold even if $D$ is not diagonal.
@@
 In another example, in the high-dimensional setting where $p\rightarrow\infty$, it is possible to identify $\Lambda^{*}\in\mathbb{R}^{p\times q}$, whose columns correspond to the top $q$ eigenvalues of $\text{Cov}(X)$.
-Under additional boundedness assumptions on the correlation matrix $D$ and the coefficient matrix $\Lambda$, one can show that there exists a matrix $O\in\mathbb{R}^{q\times q}$ such that the $\ell_{2}$ norm between each column of $\Lambda O$ and $\Lambda^{*}$ converges to zero as $p$ tends to infinity.
+Under additional boundedness assumptions on the correlation matrix $D$ and the coefficient matrix $\Lambda$, one can show that there exists a matrix $O\in\mathbb{R}^{q\times q}$ such that the $\ell_{2}$-norm between each column of $\Lambda O$ and $\Lambda^{*}$ converges to zero as $p$ tends to infinity.
 See Fan et al.
@@
 where $\Lambda_{j}$ is the $j$th element of $\Lambda\in\mathbb{R}^{p\times 1}$ and $\gamma=\mathbb{E}(Ug(U))$.
-Since there are three equations in (4) but four unknown parameters $\beta_{1},\beta_{2},\beta_{3},\gamma$, the causal parameters $\beta$ are not identifiable from these equations.
+Since there are three equations in (4) but four unknown parameters, $\beta_{1},\beta_{2},\beta_{3},\gamma$, the causal parameters $\beta$ are not identifiable from these equations.
 One possible approach to identifying $\beta$ is to assume prior knowledge about certain elements of $\beta$.
@@
 […]
-(c) Assume $||\beta||_{0}\leq 1$.
+(c) Assume $\beta||_{0}\leq 1$.
 Figure 1: Causal diagrams corresponding to models (1) and (3): $p=3,q=1.$
@@
   1. A3
-    (Sparsity) $\|\dot{\beta}\|_{0}\leq p-q-1$, where $\dot{\beta}$ refers to the true value of $\beta$.
+    (Sparsity) $\|\dot{\beta}\|_{0}\leq p-q-1$, where $\dot{\beta}$ denotes the true value of $\beta$.
 Remark 1.
-Condition A3 is significantly less restrictive compared to similar assumptions in the existing literature used for causal effect identification in this context.
+Condition A3 is significantly less restrictive than similar assumptions in the existing literature used for causal effect identification in this context.
 For example, Miao et al.
@@
 2.3 Instrumental variable
-The method of instrumental variables (IV) is a widely used approach for estimating causal relationships when unmeasured confounders exist between the exposure $X$ and the outcome $Y$.
+The method of instrumental variables is a widely used approach for estimating causal relationships when unmeasured confounders exist between the exposure $X$ and the outcome $Y$.
 Suppose we have an exogenous variable $Z$.
-For simplicity, assume that the relationships among the random variables are linear and follow the structure equation models:
+For simplicity, assume that the relationships among the random variables are linear and follow the structural equation models:
 $\begin{split}Y&=\beta X+\gamma U+\pi Z+\epsilon_{y},\\ X&=\alpha_{z}Z+\Lambda U+\epsilon_{x}.\end{split}$
 For $Z$ to be a valid instrumental variable, the following assumptions are commonly made (e.g., Wang and Tchetgen Tchetgen, 2018): $\pi=0$ (exclusion restriction), $\alpha_{z}\neq 0$ (instrumental relevance), and $\text{Cov}(U,Z)=0$ (unconfoundedness).
-Under these assumptions, one can consistently estimate $\beta$ via a two-stage least squares estimator: first, obtain the predicted exposure $\widehat{\mathbb{E}}(X\mid Z)$ by linearly regressing $X$ on $Z$, and then regress $Y$ on $\widehat{\mathbb{E}}(X\mid Z)$ to obtain an estimate of $\beta$; here $\mathbb{E}(X\mid Z)$ refers to the conditional expectation of $X$ given $Z$.
+Under these assumptions, one can consistently estimate $\beta$ via a two-stage least squares estimator: first, obtain the predicted exposure $\widehat{\mathbb{E}}(X\mid Z)$ by linearly regressing $X$ on $Z$, and then regress $Y$ on $\widehat{\mathbb{E}}(X\mid Z)$ to obtain an estimate of $\beta$.
+Here, $\mathbb{E}(X\mid Z)$ refers to the conditional expectation of $X$ given $Z$, and $\widehat{\mathbb{E}}(X\mid Z)$ refers to its estimator obtained through linear regression.
 3 Identifying causal effects via the synthetic instrument
@@
 Therefore, in general, $\widetilde{\beta}^{(3)}\neq\beta$.
-To decide among these three experts, we compare the solutions $\widetilde{\beta}^{(j)}$ and find their mode $\beta_{\text{mode}}=\mathop{\arg\max}\limits_{\beta\in\mathbb{R}^{3}}|\{j:\widetilde{\beta}^{(j)}=\beta\}|$; here $|\mathcal{S}|$ denotes the cardinality of a set $\mathcal{S}$.
+To decide among these three experts, we compare the solutions $\widetilde{\beta}^{(j)}$ and find their mode, defined as $\beta_{\text{mode}}=\mathop{\arg\max}\limits_{\beta\in\mathbb{R}^{3}}|\{j:\widetilde{\beta}^{(j)}=\beta\}|,$ where $|\mathcal{S}|$ denotes the cardinality of a set $\mathcal{S}$.
 One can easily see from Table 1 that $\beta_{\text{mode}}=\dot{\beta}$.
@@
 In total, there are $C_{p}^{q}$ different hypotheses.
-Several challenges arise when the data is moderate to high-dimensional, so that $p$ and $q$ are not small.
+Several challenges arise when the data are moderate to high-dimensional, so that $p$ and $q$ are not small.
   1. (i)
@@
 Remark 2.
-Other approaches that use the voting analogy for identification (e.g.
-Zhou et al., 2014; Guo et al., 2018) face the same challenges we present here.
+Other approaches that use the voting analogy for identification (e.g., Zhou et al., 2014; Guo et al., 2018) face the same challenges we present here.
 It is only due to the special structure of our problem that we are able to develop a method that bypasses the model selection step and addresses these challenges.
-See Sun et al.
-(2023) for another example in a different context.
 In the following, we first introduce the SIV in the context of Figure 1(b), where it is assumed that $\beta_{2}=0$.
@@
 $\begin{split}X_{1}&=\Lambda_{1}U+\epsilon_{1},\\ X_{2}&=\Lambda_{2}U+\epsilon_{2},\end{split}$ (5)
-where $\Lambda_{1}$ and $\Lambda_{2}$ are identified up to the same sign flip so that $\Lambda_{1}/\Lambda_{2}$ is identifiable.
-Eliminating $U$ from (5), we get $X_{1}-\frac{\Lambda_{1}}{\Lambda_{2}}X_{2}=\epsilon_{1}-\frac{\Lambda_{1}}{\Lambda_{2}}\epsilon_{2}$, which only depends on the error terms $\epsilon_{1}$ and $\epsilon_{2}$.
-Since $\epsilon_{2}$ is also uncorrelated with $U$, it is not difficult to see from Figure 1(b) that $SIV_{1}^{(2)}=X_{1}-\frac{\Lambda_{1}}{\Lambda_{2}}X_{2}$ satisfies the conditions for an instrumental variable for identifying $\beta_{1}$ described in Section 2.3, hence the name synthetic instrument.
+where $\Lambda_{1}$ and $\Lambda_{2}$ are identified up to the same sign flip, so that $\Lambda_{1}/\Lambda_{2}$ is identifiable.
+Eliminating $U$ from (5), we obtain $X_{1}-{\Lambda_{2}}X_{2}/{\Lambda_{1}}=\epsilon_{1}-{\Lambda_{1}}\epsilon_{2}/{\Lambda_{1}},$ which depends only on the error terms $\epsilon_{1}$ and $\epsilon_{2}$.
+Since $\epsilon_{2}$ is also uncorrelated with $U$, it is not difficult to see from Figure 1(b) that $SIV_{1}^{(2)}=X_{1}-{\Lambda_{1}}X_{2}/{\Lambda_{2}}$ satisfies the conditions for an instrumental variable for identifying $\beta_{1}$ described in Section 2.3, hence the name synthetic instrument.
 In contrast to a standard instrumental variable, the synthetic instrument is directly constructed as a linear combination of the exposures, so there is no need to measure additional exogenous variables.
-To identify $\beta_{3}$, one can similarly define $SIV_{1}^{(3)}=X_{3}-\frac{\Lambda_{1}}{\Lambda_{2}}X_{3}.$ Let $SIV^{(2)}=\left(SIV_{1}^{(2)},SIV_{3}^{(2)}\right)$.
+To identify $\beta_{3}$, one can similarly define $SIV_{3}^{(2)}=X_{3}-{\Lambda_{3}}X_{2}/{\Lambda_{2}}$.
+Let $SIV^{(2)}=\left(SIV_{1}^{(2)},SIV_{3}^{(2)}\right)$.
 One can then obtain $(\beta_{1},\beta_{3})$ using the so-called synthetic two-stage least squares:
   1. 1.
-    Fit a linear regression of $X=(X_{1},X_{2},X_{3})$ on $SIV^{(2)}=(SIV_{1}^{(2)},SIV_{3}^{(2)})$ and obtain $\widetilde{X}=E[X\mid SIV^{(2)}]$.
+    Fit a linear regression of $X=(X_{1},X_{2},X_{3})$ on $SIV^{(2)}=(SIV_{1}^{(2)},SIV_{3}^{(2)})$ and obtain $\widetilde{X}=\widehat{\mathbb{E}}[X\mid SIV^{(2)}]$ through the fitted values of the linear regression.
   2. 2.
@@
   1. 1.
-    For $j=1,2,3$, fit a linear regression of $X$ on $SIV^{(j)}$ and obtain $\widetilde{X}^{(j)}=E\left[X\mid SIV^{(j)}\right]$;
+    For $j=1,2,3$, fit a linear regression of $X$ on $SIV^{(j)}$ and obtain $\widetilde{X}^{(j)}=\widehat{\mathbb{E}}\left[X\mid SIV^{(j)}\right]$;
   2. 2.
-    Fit a linear regression of $Y$ on $\widetilde{X}^{(j)}$, fixing $\beta_{j}=0$, and obtain the coefficients $\widetilde{\beta}^{(j)}$.
+    Fit a linear regression of $Y$ on $\widetilde{X}^{(j)}$, fixing $\beta_{j}=0$, and obtain the coefficients $\widetilde{\beta}^{(j)}$;
   3. 3.
     Find the mode among $\widetilde{\beta}^{(j)},\,j=1,2,3.$
-On the surface, similar to the problems described at the beginning of Section 3.2, voting with the synthetic instrument still involves fitting three different regressions and comparing three vectors $\widetilde{\beta}^{(j)}.$ We now make two key observations regarding the properties of the synthetic instrument, which allows us to simplify Algorithm 1 into a two-stage regression procedure.
+On the surface, similar to the problems described at the beginning of Section 3.2, voting with the synthetic instrument still involves fitting three different regressions and comparing three vectors $\widetilde{\beta}^{(j)}$.
+We now make two key observations regarding the properties of the synthetic instrument, which allow us to simplify Algorithm 1 into a two-stage regression procedure.
 Observation 1.
 Let $\Lambda=(\Lambda_{1},\Lambda_{2},\Lambda_{3})$.
-For $j=1,2,3$, $SIV^{(j)}\in\mathbb{R}^{2}$ span the same linear space $\{\lambda^{T}X:\lambda\in\Lambda^{\perp}\}.$ As a result, $\widetilde{X}^{(j)}=E\left[X\mid SIV^{(j)}\right]$ does not depend on the choice of $j$ so that one only needs to run Step 1 of Algorithm 1 once.
+For $j=1,2,3$, $SIV^{(j)}\in\mathbb{R}^{2}$ span the same linear space $\{\lambda^{\mathrm{\scriptscriptstyle T}}X:\lambda\in\Lambda^{\perp}\}$.
+As a result, $\widetilde{X}^{(j)}=\mathbb{E}\left[X\mid SIV^{(j)}\right]$ does not depend on the choice of $j$, so that one only needs to run Step 1 of Algorithm 1 once.
 Observation 2.
@@
 Recall that the true value is $\dot{\beta}=\widetilde{\beta}^{(1)}=\widetilde{\beta}^{(2)}$.
-Instead of calculating $\widetilde{\beta}^{(j)},j=1,2,3$ separately for each $j$, Steps 2 and 3 in Algorithm 1 can be replaced with the following penalized regression:
-$\beta^{SIV}=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{3}}\|\ Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta\|_{2}^{2}\text{\quad subject to }\|\beta\|_{0}\leq 1,$
+Instead of calculating $\widetilde{\beta}^{(j)},\,j=1,2,3,$ separately for each $j$, Steps 2 and 3 in Algorithm 1 can be replaced with the following penalized regression:
+$\beta^{SIV}=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{3}}\|Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta\|_{2}^{2}\quad\text{subject to }\|\beta\|_{0}\leq 1,$
 where, due to Observation 1, $\widetilde{X}^{(1)}=\widetilde{X}^{(2)}=\widetilde{X}^{(3)}\equiv\widetilde{X}$.
@@
 3.4 Synthetic two-stage regularized regression
-We now formally introduce the synthetic two-stage regularized regression for the general case. Motivated by Observation 1, we introduce the following definition for the synthetic instrument.
+We now formally introduce the synthetic two-stage regularized regression for the general case. Motivated by Observation 1, we provide the following definition of the synthetic instrument.
 Definition 1 (Synthetic Instrument).
@@
 Proposition 2.
-Under models (1), (3), and Condition A2, the $SIV$ given by Definition 1 serve as valid instrumental variables for estimating the treatment effects of $X$ on $Y$.
+Under models (1), (3), and Condition A2, the SIVSIV given by Definition 1 serve as valid instrumental variables for estimating the treatment effects of $X$ on $Y$.
 To identify the causal parameter $\beta$ in the general case, we introduce the following plurality condition A4.
@@
     (Plurality rule) Let $C^{*}$ be a subset of $\{1,2,\ldots,p\}$ with cardinality $q$, and suppose that $\dot{\beta}_{C^{*}}\neq 0$.
-The synthetic two-stage least squares coefficient obtained by assuming $\beta_{C^{*}}=0$ is given by $\widetilde{\beta}^{C^{*}}=\underset{\beta\in\mathbb{R}^{p}:\beta_{C^{*}}=0}{\operatorname*{arg\,min}}\,\mathbb{E}(Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta)^{2}$, where $\widetilde{X}=\mathbb{E}(X\mid SIV)$.
+The synthetic two-stage least squares coefficient obtained by assuming $\beta_{C^{*}}=0$ is given by $\widetilde{\beta}^{C^{*}}=\underset{\beta\in\mathbb{R}^{p}:\beta_{C^{*}}=0}{\operatorname*{arg\,min}}\,\mathbb{E}\bigl(Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta\bigr)^{2},$ where $\widetilde{X}=\widehat{\mathbb{E}}(X\mid SIV)$.
 The plurality rule assumes that $\max\limits_{\beta\in\mathbb{R}^{p}}\,|\{C^{*}:\widetilde{\beta}^{C^{*}}=\beta\}|\leq q.$
@@
 In general, there are $C_{p-s}^{q}$ experts making correct hypotheses.
-If $s<p-q$, then there are at least $q+1$ experts making correct hypotheses.
+If s<pqs<p-q, then there are at least $q+1$ experts making correct hypotheses.
 The plurality rule assumes that no more than $q$ incorrect hypotheses lead to the same synthetic two-stage least squares coefficient.
-This assumption is similar in spirit to the plurality assumption used in the invalid IV literature (e.g.
-Guo et al., 2018).
-We discussed this assumption in more detail in Section LABEL:sec:disA4 of the supplementary material.
+This assumption is similar in spirit to the plurality assumption used in the invalid IV literature (e.g., Guo et al., 2018).
+We further discuss Assumption A4 in Section LABEL:sec:disA4 of the supplementary material, where we argue that its violation is unlikely.
 In parallel to Observation 2, we have the following theorem.
 Theorem 1 (Synthetic two-stage regularized regression).
-Suppose that models (1), (3) and conditions A1, A2, and A4 hold.
+Suppose that models (1), (3), and Conditions A1, A2, and A4 hold.
   1. 1.
@@
     If A3 fails, then $\dot{\beta}$ is not identifiable, and for all $\widetilde{\beta}\in\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;\mathbb{E}(Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta)^{2}$, we have $\|\widetilde{\beta}\|_{0}\geq p-q$.
-An important feature of Theorem 1 is that given $q$, it is possible to test the sparsity condition A3 from the observed data.
+An important feature of Theorem 1 is that, given $q$, it is possible to test the sparsity condition A3 from the observed data.
 In particular, it shows that under models (1), (3), Conditions A1, A2, and the plurality rule A4, the following three statements are equivalent:
@@
   3. (3)
-    The most sparse least-squares solution to the second-stage regression has an $\ell_{0}$-norm smaller than $p-q$, i.e.,
+    The most sparse least-squares solution to the second-stage regression has an $\ell_{0}$-norm smaller than pqp-q, i.e.,
     $\min\limits_{\widetilde{\beta}\in\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;\mathbb{E}(Y-\widetilde{X}^{\mathrm{\scriptscriptstyle T}}\beta)^{2}}\|\widetilde{\beta}\|_{0}<p-q.$ (6)
-It is worth noting that (6) can be checked from the observed data distribution, so that one may develop a consistent test for Condition A3 and identifiability of $\beta$ under models (1), (3) and conditions A1, A2, and A4.
+It is worth noting that (6) can be checked from the observed data distribution, so that one may develop a consistent test for Condition A3 and the identifiability of $\beta$ under models (1), (3), and Conditions A1, A2, and A4.
 See Algorithm 2 below for more details.
@@
 $\begin{split}\widehat{\bf X}&=\widehat{\mathbb{E}}({\bf X}\mid\widehat{SIV}),\\ \widehat{\beta}&=\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}{\|{\bf Y}-\widehat{\bf X}\beta\|_{2}^{2}}\quad\text{subject to }\|\beta\|_{0}\leq k,\end{split}$ (7)
-where $k$ is a tuning parameter, $\widehat{\Lambda}$ is an estimator of the loading matrix, and $\widehat{SIV}=\bm{X}B_{\widehat{\Lambda}^{\perp}}.$
-Several different estimators have been proposed for determining the number of latent factors in a factor model.
-In our simulations and data analysis, we will use the estimator developed by Onatski (2010) to obtain $\widehat{q}$, as it is applicable in both low- and high-dimensional settings.
-Similarly, numerous estimators have been developed for the loading matrix $\Lambda$.
-For low-dimensional settings where $p$ is fixed, we recommend using the maximum likelihood estimator for $\Lambda$ by optimizing the log-likelihood under the assumption of multivariate normality.
-For high-dimensional settings, we suggest estimating $\Lambda$ using the principal component analysis (PCA) method (Bai, 2003), which provides a consistent estimator for the loading matrix without assuming the covariance matrix $\text{Cov}(\epsilon_{X})$ to be diagonal.
-Finally, we use cross-validation to select the tuning parameter $k$.
-Algorithm 2 summarizes our estimation procedure.
+where $k$ is a tuning parameter, $\widehat{\Lambda}$ is an estimator of the loading matrix, and $\widehat{SIV}={\bf X}B_{\widehat{\Lambda}^{\perp}}.$
+Several estimators have been proposed to determine the number of latent factors in a factor model.
+In our simulations and data analysis, we use the estimator developed by Onatski (2010) to obtain $\widehat{q}$, as it is applicable in both low- and high-dimensional settings.
+Likewise, various methods exist for estimating the loading matrix $\Lambda$.
+In low-dimensional settings where $p$ is fixed, we recommend the maximum likelihood estimator of $\Lambda$, obtained by maximizing the log-likelihood under multivariate normality.
+In high-dimensional settings, we estimate $\Lambda$ using principal component analysis (PCA) (Bai, 2003), which yields a row-consistent estimator for $\Lambda$, that is, each estimated row $\widehat{\Lambda}_{i\cdot}$ consistently estimates its true counterpart ${\Lambda}_{i\cdot}$.
+This approach does not require the covariance matrix $\text{Cov}(\epsilon_{X})$ to be diagonal.
+Finally, we use cross-validation to select the tuning parameter $k$. Algorithm 2 summarizes our estimation procedure.
 Algorithm 2  The synthetic two-stage regularized regression
 Input: ${\bf X}\in\mathbb{R}^{n\times p}$ (centered), ${\bf Y}\in\mathbb{R}^{n\times 1}$
-1:Obtain $\widehat{q}$ from $\bm{X}$ (e.g., Onatski, 2010).
+1:Obtain $\widehat{q}$ from $\bf X$} (e.g., Onatski, 2010).
 2:if $n>p$ then obtain $\widehat{\Lambda}\in\mathbb{R}^{p\times q}$ via maximum likelihood estimation, assuming multivariate normality;
@@
 This can be obtained, for example, using the Null function from the MASS package in R.
-5:Obtain $\widehat{SIV}=\bm{X}B_{\widehat{\Lambda}^{\perp}}$.
-6:Obtain $\widehat{\bf X}=\widehat{\mathbb{E}}({\bf{X}}\mid\widehat{SIV})$ via ordinary least squares.
+5:Obtain $\widehat{SIV}={\bf X}B_{\widehat{\Lambda}^{\perp}}$.
+[…]
 7:Obtain $\widehat{\beta}$ via (7), where the tuning parameter $k$ is selected via 10-fold cross-validation.
 8:if $\widehat{q}+\widehat{k}<p$ then output $\widehat{\beta}$;
-9:else $\beta$ is not identifiable.
+[…]
 4.2 Theoretical properties
 In this section, we study the theoretical properties of the estimator $\widehat{\beta}$ in Algorithm 2.
-We consider two paradigms: (1) low-dimensional settings where the dimension of exposure, $p$, is fixed; and (2) high-dimensional settings where $p$ grows with the sample size $n$.
+We consider two paradigms: (1) low-dimensional settings, where the dimension of exposure $p$ is fixed; and (2) high-dimensional settings, where $p$ grows with the sample size $n$.
 For the former, we show that under mild regularity conditions, $\widehat{\beta}$ is $\sqrt{n}$-consistent.
@@
 We also demonstrate variable selection consistency in both scenarios.
-In our theoretical results, we do not require $\widehat{\Lambda}=\Lambda$.
-We only need certain norms of $\widehat{\Lambda}$ to be consistent with $\Lambda$, which can be achieved by classical estimators.
+In our theoretical results, we do not require $\widehat{\Lambda}=\Lambda$; we only need certain norms of $\widehat{\Lambda}$ to be consistent with those of $\Lambda$, which can be achieved by classical estimators.
 We first introduce assumptions for the low-dimensional case.
@@
   1. B1
-    All coefficients ${\Lambda}$, $\beta$, and the function $g(\cdot)$ in models (1) and (3) are fixed and do not change as $n$ goes to infinity.
+    All coefficients $\Lambda$}, $\beta$, and the function $g(\cdot)$ in models (1) and (3) are fixed and do not change as $n\to\infty$.
   2. B2
     $U_{i}$, $\epsilon_{x,i}$, and $\epsilon_{y,i}$ are independent random draws from the joint distribution of $(U,\epsilon_{x},\epsilon_{y})$ such that $E(\epsilon_{x})=\bm{0}$, $E(U)=\bm{0}$, $\text{Cov}(\epsilon_{x})=D$, $\text{Cov}(U)=I_{q}$, and $(U,\epsilon_{x},\epsilon_{y})$ are mutually independent.
-Furthermore, assume that $\text{Var}({\epsilon}_{y})={\sigma}^{2}$ and $\max_{1\leq j\leq p}\text{Var}(X_{j})=\sigma_{x}^{2}$; these parameters are fixed and do not change as $n$ goes to infinity.
+Furthermore, assume that $\text{Var}(\epsilon_{y})=\sigma^{2}$ and $\max_{1\leq j\leq p}\text{Var}(X_{j})=\sigma_{x}^{2}$; these parameters are fixed and do not change as $n\to\infty$.
   3. B3
@@
   4. B4
-    Let ${\Sigma}_{\widetilde{X}}=\text{Cov}(\widetilde{X})$.
-We assume $\min_{\theta\in\mathbb{R}^{p},0<\|\theta\|_{0}\leq 2s}\frac{\theta^{\mathrm{\scriptscriptstyle T}}\Sigma_{\widetilde{X}}\theta}{\|\theta\|_{2}^{2}}>c$ for some positive constant $c$.
+[…]
 Conditions B1B2 are standard assumptions for the low-dimensional setting.
-Given Condition A2, Condition B3 assumes that the estimator for factor loading is root-$n$ consistent.
+Given Condition A2, Condition B3 requires that the estimator for factor loadings is root-$n$ consistent.
 Condition B4 is the population version of the sparse eigenvalue condition (Raskutti et al., 2011, Assumption 3(b)).
@@
 Then $\mathbb{P}(\widehat{\mathcal{A}}=\mathcal{A})\rightarrow 1$ as $n\rightarrow\infty.$
-In Theorem 2, it is assumed that $\widehat{k}=s$. This is a standard condition in the $\ell_{0}$-optimization literature (e.g. Raskutti et al., 2011; Shen et al., 2013).
+In Theorem 2, it is assumed that $\widehat{k}=s$. This is a standard condition in the $\ell_{0}$-optimization literature (e.g., Raskutti et al., 2011; Shen et al., 2013).
 Next, we consider the high-dimensional case and demonstrate that our estimator exhibits properties similar to those of standard regularized estimators in the high-dimensional statistics literature, including a Lasso-type error bound and consistency in variable selection.
@@
   1. C1
-    $sq^{2}\log(p)\log(n)/n\rightarrow 0$, $n=O(p)$, and $q+\log(p)\lesssim\sqrt{n}$, where $x\lesssim y$ means there exist a constant $C$ such that $x\leq Cy$.
+    $sq^{2}\log(p)\log(n)/n\rightarrow 0$, $n=O(p)$, and $q+\log(p)\lesssim\sqrt{n}$, where $x\lesssim y$ means there exists a constant $C$ such that $x\leq Cy$.
   2. C2
@@
 We also assume $\epsilon_{y}$ is independent of $(X,U)$ and $\epsilon_{x}$ is independent of $U$.
-Furthermore, assume ${\epsilon}_{y}$, $\epsilon_{x,j}$, and $X_{j}$ are sub-Gaussian random variables with sub-Gaussian parameters ${\sigma}^{2}$, $\widetilde{\sigma}_{j}^{2}$, and $\sigma_{j}^{2}$.
+Furthermore, assume $\epsilon_{y}$, $\epsilon_{x,j}$, and $X_{j}$ are sub-Gaussian random variables with sub-Gaussian parameters $\sigma^{2}$, $\widetilde{\sigma}_{j}^{2}$, and $\sigma_{j}^{2}$, respectively.
 The parameters satisfy $\sigma^{2}\leq C_{4}$, and $C_{5}\leq\widetilde{\sigma}_{j}^{2},\sigma_{j}^{2}\leq C_{6}$ for some constants $C_{4},C_{5},C_{6}>0$.
@@
 Condition C2 is a standard assumption in high-dimensional factor analysis (Fan et al., 2013; Shen et al., 2016) for loading identification.
-Condition C3 assumes that the exposures $X_{j}$ are sub-Gaussian, and the noise level is bounded.
+Condition C3 assumes that the exposures $X_{j}$ are sub-Gaussian and that the noise level is bounded.
 Condition C4 is a standard assumption on minimum signal strength.
 Theorem 3.
-Assuming that Conditions A1A4 and C1C3 hold, and the tuning parameter $\widehat{k}=s$, we have:
+Assume that Conditions A1A4 and C1C3 hold, and that the tuning parameter satisfies $\widehat{k}=s$. Then:
   1. 1.
@@
 where $b$ is the best linear predictor of $U$ from $X$.
-Their result depends on (i) $\|b\|_{2}=O_{p}\left(\frac{1}{\sqrt{p}}\right)$ and (ii) the term $(U^{\mathrm{\scriptscriptstyle T}}\gamma-X^{\mathrm{\scriptscriptstyle T}}b)+\epsilon_{y}$ being independent of $X$ under their joint Gaussian assumption.
-However, these assumptions may not hold in our model (3), where $g(U)$ is an unknown function.
+Their result depends on (i) $\|b\|_{2}=O\left({1}/{\sqrt{p}}\right)$ and (ii) the term $(U^{\mathrm{\scriptscriptstyle T}}\gamma-X^{\mathrm{\scriptscriptstyle T}}b)+\epsilon_{y}$ being independent of $X$ under their joint Gaussian assumption.
+In general, these conditions fail to hold under model (3), where $g(U)$ is an unknown function.
 4.3 Extension to nonlinear settings
@@
 Revisiting model (3), we remove the assumption of linearity, allowing both the treatment $X$ and the unmeasured confounder $U$ to influence the outcome $Y$ through nonlinear relationships:
-$Y=f(X;\beta)+g(U)+\epsilon_{y}.$
+$Y=f(X;\beta)+g(U)+\epsilon_{y}$ (8)
 In this model, the treatment influences the outcome through the nonlinear causal function $f(\cdot;\beta)$, with $\beta\in\mathbb{R}^{p}$ as the parameter of interest.
@@
 $\mathbb{E}\{SIV(Y-f(X;\beta))\}=\mathbb{E}\{SIV(g(U)+\epsilon_{y})\}=0,$
-assuming $\beta$ is set to its true value. Following this, we define the population GMM loss:
+when $\beta$ is set to its true value. Following this, we define the population GMM loss:
 $G(\beta)=\left\|\mathbb{E}\{SIV(Y-f(X;\beta))\}\right\|_{2}^{2}.$ (9)
-For the function $f(X;\beta)$, let $\partial f(X;\beta)/\partial\beta$ denote the $p\times 1$ column vector $\left({\partial f(X;\beta)}/{\partial\beta_{1}},\ldots,{\partial f(X;\beta)}/{\partial\beta_{p}}\right)^{\mathrm{\scriptscriptstyle T}}$ and let $\partial f(X;\beta)/\partial\beta^{\mathrm{\scriptscriptstyle T}}$ denote the $1\times p$ row vector $\left({\partial f(X;\beta)}/{\partial\beta_{1}},\ldots,{\partial f(X;\beta)}/{\partial\beta_{p}}\right)$.
+For the function $f(X;\beta)$, let $\partial f(X;\beta)/\partial\beta$ denote the $p\times 1$ column vector $({\partial f(X;\beta)}/{\partial\beta_{1}}$ $\ldots,{\partial f(X;\beta)}/{\partial\beta_{p}})^{\mathrm{\scriptscriptstyle T}},$ and let $\partial f(X;\beta)/\partial\beta^{\mathrm{\scriptscriptstyle T}}$ denote the $1\times p$ row vector $({\partial f(X;\beta)}/{\partial\beta_{1}},\ldots,$ ${\partial f(X;\beta)}/{\partial\beta_{p}})$.
 We make the following assumptions, denoted as D1 and D2, which are generalizations of Conditions A1 and A4 in the nonlinear setting.
@@
   2. D2
-    (Nonlinear plurality rule) Let $C^{*}$ be a subset of $\{1,2,\ldots,p\}$ with cardinality $q$ and $\dot{\beta}_{C^{*}}\neq 0$, and let the synthetic GMM obtained by assuming $\beta_{{C^{*}}}=0$ be $\widetilde{\beta}^{C^{*}}=\underset{\beta\in\mathbb{R}^{p}:\beta_{C^{*}}=0}{\operatorname*{arg\,min}}G(\beta).$ The plurality rule assumes that $\widetilde{\beta}^{{C^{*}}}$ is uniquely defined and that $\max\limits_{\beta\in\mathbb{R}^{p}}|\{C^{*}:\widetilde{\beta}^{C^{*}}=\beta\}|\leq q.$
+    (Nonlinear plurality rule) Let $C^{*}$ be a subset of $\{1,2,\ldots,p\}$ with cardinality $q$ and $\dot{\beta}_{C^{*}}\neq 0$, and let the synthetic GMM estimator obtained by assuming $\beta_{C^{*}}=0$ be $\widetilde{\beta}^{C^{*}}=\underset{\beta\in\mathbb{R}^{p}:\beta_{C^{*}}=0}{\operatorname*{arg\,min}}G(\beta).$ The plurality rule assumes that $\widetilde{\beta}^{C^{*}}$ is uniquely defined and that $\max\limits_{\beta\in\mathbb{R}^{p}}|\{C^{*}:\widetilde{\beta}^{C^{*}}=\beta\}|\leq q.$
 The following theorem states that $\beta$ is identifiable using synthetic GMM in a manner parallel to Theorem 1, but within a nonlinear setting.
 Theorem 4 (Synthetic Generalized Method of Moments).
-Suppose that models (1), (3) and conditions A2, D1, and D2 hold.
+Suppose that models (1), (3), and Conditions A2, D1, and D2 hold.
   1. 1.
-    If A3 holds, then $\dot{\beta}$ is identifiable via $\dot{\beta}=\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;G(\beta)$, subject to $||\beta||_{0}<p-q$.
+    If A3 holds, then $\dot{\beta}$ is identifiable via $\dot{\beta}=\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;G(\beta)$, subject to $\|\beta\|_{0}<p-q$.
   2. 2.
-    If A3 fails, then $\dot{\beta}$ is not identifiable, and for all $\widetilde{\beta}\in\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;G(\beta)$, we have $||\widetilde{\beta}||_{0}\geq p-q$.
-We discuss how to estimate $\beta$ in the finite sample setting. We replace the expectation in equation (8) with the empirical average. Specifically, we consider the following loss function:
-$G_{n}(\beta)=\left(\frac{1}{n}\sum^{n}_{i=1}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right)^{\mathrm{\scriptscriptstyle T}}W\left(\frac{1}{n}\sum^{n}_{i=1}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right),$
+    If A3 fails, then $\dot{\beta}$ is not identifiable, and for all $\widetilde{\beta}\in\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;G(\beta)$, we have $\|\widetilde{\beta}\|_{0}\geq p-q$.
+We now discuss the estimation of $\beta$ in finite samples.
+A natural approach is to replace the expectation in equation (9) with its empirical counterpart.
+Specifically, we consider the following loss function:
+$G_{n}(\beta)=\left(\frac{1}{n}\sum_{i=1}^{n}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right)^{\mathrm{\scriptscriptstyle T}}W\left(\frac{1}{n}\sum_{i=1}^{n}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right),$
 where $W\in\mathbb{R}^{(p-q)\times(p-q)}$ is a weight matrix.
-The oracle weight matrix that achieves the highest efficiency in GMM is the inverse of $\mathbb{E}\left[\text{SIV}(Y-f(X;\dot{\beta}))^{2}\text{SIV}^{\mathrm{\scriptscriptstyle T}}\right]=\widetilde{\sigma}^{2}\text{Cov}(\text{SIV})$, where $\widetilde{\sigma}^{2}=\text{Var}(g(U)+\epsilon_{y})$ (Burgess et al., 2017).
-In practice, we estimate the weight matrix $W$ using the inverse of empirical covariance: $\widehat{\text{Cov}}^{-1}(\widehat{SIV})$.
-Finally, let $\bm{X}\in\mathbb{R}^{n\times p}$, $\bm{Y}\in\mathbb{R}^{n\times 1}$, $f(\bm{X};\beta)\in\mathbb{R}^{n\times 1}$, and $\bm{SIV}\in\mathbb{R}^{n\times(p-q)}$ denote the relevant random matrices in the finite sample setting.
-We consider the following optimization problem:
+We now discuss the estimation of $\beta$ in finite samples.
+A natural approach is to replace the expectation in equation (9) with its empirical counterpart.
+Specifically, we consider the following loss function:
+$G_{n}(\beta)=\left(\frac{1}{n}\sum_{i=1}^{n}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right)^{\mathrm{\scriptscriptstyle T}}W\left(\frac{1}{n}\sum_{i=1}^{n}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right),$
+[…]
+Finally, let $\bm{X}\in\mathbb{R}^{n\times p}$, $\bm{Y}\in\mathbb{R}^{n\times 1}$, $f(\bm{X};\beta)\in\mathbb{R}^{n\times 1}$, and $\bm{SIV}\in\mathbb{R}^{n\times(p-q)}$ denote the relevant random matrices in the finite-sample setting.
+In the low-dimensional setting, where the number of instruments pqp-q is fixed and smaller than $n$, we consider the following optimization problem:
 $\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\left\|\bm{SIV}({\bm{SIV}^{\mathrm{\scriptscriptstyle T}}\bm{SIV}})^{-1}{\bm{SIV}^{\mathrm{\scriptscriptstyle T}}(\bm{Y}-f(\bm{X};\beta))}\right\|^{2}_{2}\quad\text{subject to }\|\beta\|_{0}\leq k.$ (10)
 The unconstrained optimization in (10) is also referred to as the nonlinear two-stage least squares estimator in the literature (Amemiya, 1974).
-Under the linear setting where $f(\bm{X};\beta)=\bm{X}\beta$, (9) has a similar form to (7).
+Under the linear setting where $f(\bm{X};\beta)=\bm{X}\beta$, equation (10) has a similar form to (7).
 Optimization (10) can be accelerated using a splicing approach, which is designed to efficiently solve the best subset selection problem (Zhu et al., 2020; Zhang et al., 2023).
+We establish the theoretical properties of (10) in Section LABEL:sec:appthmnonlinear of the supplementary material, showing that under mild regularity conditions, the proposed estimator achieves both consistency and variable selection consistency in the low-dimensional setting.
+Remark 4.
+In high-dimensional settings where the dimension of $\operatorname{SIV}$, namely pqp-q, exceeds the sample size $n$, the matrix $\operatorname{SIV}^{\mathrm{\scriptscriptstyle T}}\operatorname{SIV}$ is not invertible.
+In this case, one may follow the approach of Belloni et al.
+(2018, p.
+35) and choose the weight matrix $W$ as
+$W=\operatorname{diag}\!\left(\operatorname{Var}\!\left(\mathbb{E}_{n}\!\left[\operatorname{SIV}(Y-f(X;\widetilde{\beta}))\right]\right)\right)^{-1},$
+where $\widetilde{\beta}$ denotes an initial estimator of $\beta$.
 5 Simulation studies
@@
 Next, we evaluate the algorithm in the context of a nonlinear outcome model in Section 5.2.
-In Section LABEL:sec:MS of the supplementary material, we explore an alternative moment selection estimator inspired by the work of Andrews (1999) and compare it with our proposed estimator.
+We provide additional simulation results in the supplementary material.
+In Section LABEL:sec:simuweakeffect, we present simulation results for various estimators under dense confounding with many weak effects.
+In Section LABEL:sec:MS, we explore an alternative moment selection estimator, inspired by the work of Andrews (1999), and compare it with our proposed estimator.
+Section LABEL:sec:CI_sup discusses the construction of confidence intervals for selected causal variables using the ivreg function.
+Finally, in Section LABEL:sec:countdata, we extend the SIV algorithm to confounded count data, where the effect of the unmeasured confounder on the outcome cannot be additively separated from the effect of the treatment.
 5.1 Simulation studies with a linear outcome model
@@
 In our simulations, we let $q=3$, $s=5$, and $\beta=(1,1,1,1,1,0,\ldots,0)^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{p}$.
-Each element in $\Lambda_{j,k}$ and $\gamma_{k}$ is independently generated from $\text{Uniform}(-1,1)$ for $j=1,\ldots,p$, and $k=1,\ldots,q$.
+Each element in $\Lambda_{j,k}$ and $\gamma_{k}$ is independently generated from $\text{Uniform}(-1,1)$ for $j=1,\ldots,p$ and $k=1,\ldots,q$.
 The hidden variables $U_{i,k}$ follow i.i.d.
 standard normal distributions for $i=1,\ldots,n$ and $k=1,\ldots,q$.
-The random errors are generated as $\epsilon_{x}\sim\mathbb{N}(0,\sigma^{2}_{x}I_{p})$ and $\epsilon_{y}\sim\mathbb{N}(0,\sigma^{2})$, where $\sigma_{x}=2$ and $\sigma=5$.
+The random errors are generated as $\epsilon_{x}\sim\mathbb{N}(0,\sigma_{x}^{2}I_{p})$ and $\epsilon_{y}\sim\mathbb{N}(0,\sigma^{2})$, where $\sigma_{x}=2$ and $\sigma=5$.
 We evaluate the performance of our method under the following two settings: (i) low-dimensional cases: $p=100$ and $n\in\{200,600,1000,\ldots,5000\}$; (ii) high-dimensional cases: $n=500$ and $p\in\{500,750,1000,\ldots,3000\}$.
 All simulation results are based on $1000$ Monte Carlo runs.
-The data-generating mechanism is designed to mimic certain features of the real application in Section 6; see Section LABEL:sec:comparison_sim&data in the supplementary material for more details.
+The data-generating mechanism is designed to mimic key features of the real application in Section 6; see Section LABEL:sec:comparison_sim&data of the supplementary material for a comparison between the real and synthetic data.
 We compare the following methods in our simulations.
@@
   3. 3.
-    (Trim, Ćevid et al., 2020): We implement Ćevid et al. (2020)’s method using the code available from https://github.com/zijguo/Doubly-Debiased-Lasso.
+    (Trim, Ćevid et al., 2020): We implement Ćevid et al.
+(2020)’s method using the code available from https://github.com/zijguo/Doubly-Debiased-Lasso, with an update detailed in Section LABEL:sec:update_supp of the supplementary material.
   4. 4.
@@
 (2023)’s method using the code available from https://www.tandfonline.com/doi/suppl/10.1080/01621459.2021.2023551.
-For the high-dimensional settings, their method cannot be applied directly because $\xi$ in their method cannot be solved by ordinary least squares.
-Therefore, we replace ordinary least squares with the lasso, where the tuning parameter is selected by 10-fold cross-validation.
+For the high-dimensional settings, their method cannot be applied directly because $\xi$ in their procedure cannot be solved by ordinary least squares.
+Therefore, we replace ordinary least squares with the Lasso, with the tuning parameter selected by 10-fold cross-validation.
   5. 5.
-    (IV-Lasso): Motivated by a reviewer’s comment, we consider the following “IV-Lasso” estimator.
-In the first step, we use Lasso to select active predictors and define $\widehat{\mathcal{A}}=\{j:\widetilde{\beta}_{j}\neq 0\}$, where $\widetilde{\beta}$ is the Lasso estimator described above.
-In the next step, we use $\widehat{SIV}$ as instrumental variables to estimate the treatment effects of $X_{\widehat{\mathcal{A}}}$ on $Y$ using the ivreg package, denoted as $\widehat{\beta}^{\;\text{IV-Lasso}}_{\widehat{\mathcal{A}}}$, and set $\widehat{\beta}^{\;\text{IV-Lasso}}_{\widehat{\mathcal{A}}^{c}}=0$.
+    (IV-Lasso) Motivated by a reviewer’s suggestion, we consider the following two-step “IV-Lasso” procedure.
+First, we apply the Lasso as a screening step to identify candidate causal predictors of $Y$.
+Specifically, we solve
+    $\widetilde{\beta}=\underset{\beta\in\mathbb{R}^{p}}{\arg\min}\left\{\|{\bf Y}-\widehat{{\bf X}}\beta\|_{2}^{2}+\lambda\|\beta\|_{1}\right\},$
+[…]
 (a) $p=100$, and $n$ varies from $200$ to $5000$.
@@
 In low-dimensional cases, as illustrated in Figure 2(a), the bias of the Lasso estimator stabilizes as the sample size grows.
-The Lasso estimator does not account for unmeasured confounding and is, therefore, not expected to perform well.
+The Lasso estimator does not account for unmeasured confounding and is therefore not expected to perform well.
 The Lasso, Trim, and Null methods exhibit considerable $\ell_{1}$-estimation bias relative to other methods.
@@
 This advantage arises from the $\ell_{0}$-optimization employed in the SIV method, which enables more accurate variable selection, as illustrated in Figure 3(a) later.
+To ensure a fair comparison across all methods, we use cross-validation to select $\lambda$.
+However, it is well known that cross-validation can lead to overfitting and an increased false discovery rate.
+In practice, researchers may adopt the “one-standard-error” rule, which selects the largest $\lambda$ such that the cross-validation error remains within one standard error of its minimum (Hastie et al., 2009; Kang et al., 2016).
+In the simulation settings we have considered, we find that the “one-standard-error rule” allows the IV-Lasso method to perform as well as the SIV method.
+See Section LABEL:sec:ivlasso1se of the supplementary material for details on how the “one-standard-error” rule improves the empirical performance of IV-Lasso.
 For the high-dimensional settings, we can see from Figure 2(b) that our SIV method consistently outperforms the comparison methods.
@@
 The Trim method, designed specifically for high-dimensional settings, outperforms both the Lasso and Null methods.
-However, our estimator still shows a much smaller bias compared to the Trim estimator.
+However, our estimator still shows a much smaller bias than the Trim estimator.
+In Section LABEL:sec:trim_sup of the supplementary material, we provide additional discussion, simulations, and comments on why the Trim estimator performs less favorably compared to our estimator.
 We also observe from Figure 2(b) that the IV-Lasso estimator underperforms compared to the naive Lasso estimator in high-dimensional settings.
-This underperformance occurs because the naive Lasso method applies $\ell_{1}$ penalization, which drives the coefficients of incorrectly selected $\widehat{\beta}_{i}$ towards zero.
+This underperformance occurs because the naive Lasso method applies $\ell_{1}$-penalization, which drives the coefficients of incorrectly selected $\widehat{\beta}_{i}$ towards zero.
 In contrast, the second step of IV-Lasso negates this shrinkage, causing the coefficients of incorrectly selected $\widehat{\beta}_{i}$ to diverge from zero, thereby increasing estimation bias.
-[…]
+Since the underlying $\beta$ is sparse, we also report the performance of variable selection for all the methods in Figure 3.
+All these methods correctly classify the true causes of the outcome as active exposures, that is, $\widehat{\mathcal{A}}\supseteq\mathcal{A}$.
+Thus, we only report the average false discovery rates, denoted as $\{\widehat{\mathcal{A}}\setminus\mathcal{A}\}/\#\widehat{\mathcal{A}}$, across 1000 Monte Carlo runs.
+It is evident that our proposal achieves the lowest false discovery rate among all the methods in both the low- and high-dimensional settings.
 We further evaluate the performance of our proposed algorithm in settings with non-diagonal $\text{Cov}(\epsilon_{x})$.
@@
 For the low-dimensional setting, we randomly select 20 pairs of $i,j\in\{1,2,\ldots,p\}$ and set $D_{i,j}=D_{j,i}=1$.
-The list of pairs is provided in Section LABEL:sec:detail:nondiagonal of supplementary material.
-In the high-dimensional settings, we set $D_{i,j}=4\times 0.3^{|i-j|}$.
+The list of pairs is provided in Section LABEL:sec:detail:nondiagonal of the supplementary material.
+In high-dimensional settings, we set $D_{i,j}=4\times 0.3^{|i-j|}$.
 All other aspects of the data-generating mechanism remain unchanged.
@@
 These are implemented using the R packages rpca and prcomp, respectively.
-The simulation results, presented in Figures 4 and 5
-illustrate the $\ell_{1}$-estimation errors and false discovery rates for various methods.
+The simulation results, presented in Figures 4 and 5, illustrate the $\ell_{1}$-estimation errors and false discovery rates for various methods.
 These methods exhibit similar trends in $\ell_{1}$-error performance as observed in the previous setting with uncorrelated errors.
@@
 standard normal distributions for $i=1,\ldots,n$ and $k=1,\ldots,q$.
-The random errors are generated by $\epsilon_{x}\sim\mathbb{N}(0,\sigma^{2}_{x}I_{p})$ and $\epsilon_{y}\sim\mathbb{N}(0,\sigma^{2})$, where $\sigma_{x}=2$ and $\sigma=1$.
+The random errors are generated as $\epsilon_{x}\sim\mathbb{N}(0,\sigma^{2}_{x}I_{p})$ and $\epsilon_{y}\sim\mathbb{N}(0,\sigma^{2})$, where $\sigma_{x}=2$ and $\sigma=1$.
 We evaluate the performance of our estimators with $n\in\{1000,\ldots,5000\}$.
 All simulation results are based on 1000 Monte Carlo runs.
-For comparison, we did not implement the other methods in Section 5.1, as they are not designed for non-linear outcome models.
+For comparison, we did not implement the other methods in Section 5.1, as they are not designed for nonlinear outcome models.
 Instead, we consider a popular method for addressing endogeneity in high-dimensional settings (e.g.
@@
     (U-hat1): First, we obtain $\widehat{\bm{U}}\in\mathbb{R}^{n\times q}$ using the equation $\widehat{\bm{U}}=\bm{X}\widehat{\text{Cov}}(X)^{-1}\widehat{\Lambda}$.
-Next, we obtain $\widehat{\beta}$ by solving the following optimization problem: $\widehat{\beta}=\underset{\beta\in\mathbb{R}^{p},\gamma\in\mathbb{R}^{q}}{\operatorname*{arg\,min}}\|\bm{Y}-f(\bm{X};\beta)-\widehat{\bm{U}}\gamma\|_{2}^{2}\quad\text{subject to}\quad\|\beta\|_{0}\leq k.$
+Next, we obtain $\widehat{\beta}$ by solving the following optimization problem:
+[…]
   3. 3.
-    (U-hat2): We first obtain $\widehat{\bm{U}}\in\mathbb{R}^{n\times q}$ similarly.
-We then obtain $\widehat{\beta}$ by solving the following optimization problem: $\widehat{\beta}=\underset{\beta\in\mathbb{R}^{p},\gamma\in\mathbb{R}^{q}}{\operatorname*{arg\,min}}\|\bm{Y}-f(\bm{X};\beta)-\widehat{\bm{U}}^{3}\gamma\|_{2}^{2}\quad\text{subject to}\quad\|\beta\|_{0}\leq k,$ where $\widehat{\bm{U}}^{3}\in\mathbb{R}^{n\times q}$ is defined as $\{\widehat{\bm{U}}^{3}\}_{i,j}=\{\widehat{\bm{U}}_{i,j}\}^{3}$.
-Note this method assumes knowledge of the specific form of $g(U)$, which is generally not available in practice.
-Figure 6 graphically presents the $\ell_{1}$-estimation error for all estimators.
-The results show that U-hat1 and U-hat2 have similar performance, with bias stabilizing as the sample size increases.
-In contrast, the bias of our estimator decreases with larger samples, indicating that our method consistently estimates the causal parameter $\beta$ in nonlinear outcome models.
+    (U-hat2): We first obtain $\widehat{\bm{U}}\in\mathbb{R}^{n\times q}$ similarly. We then obtain $\widehat{\beta}$ by solving the following optimization problem:
+[…]
+    where $\widehat{\bm{U}}^{3}\in\mathbb{R}^{n\times q}$ is defined as $\{\widehat{\bm{U}}^{3}\}_{i,j}=\{\widehat{\bm{U}}_{i,j}\}^{3}$.
+Note that this method assumes knowledge of the specific form of $g(U)$, which is generally not available in practice.
+Figure 6 displays the $\ell_{1}$-estimation errors for all estimators.
+The results indicate that U-hat1 and U-hat2 perform similarly, with their biases stabilizing as the sample size increases.
+In contrast, the bias of the proposed estimator decreases with larger sample sizes, demonstrating its consistency in estimating the causal parameter $\beta$ under nonlinear outcome models.
+In Section LABEL:sec:nonlinear,nondigonal of the supplementary material, we further consider a setting with a nonlinear outcome model and a non-diagonal $\text{Cov}(\epsilon_{x})$.
+The results suggest that the SIV method remains consistent in this complex setting.
 Remark 5.
@@
 While the $U$-hat method has been proposed in the literature (e.g., Wang and Blei, 2019), its validity has been widely challenged (e.g., Ogburn et al., 2019; Grimmer et al., 2023).
-From this perspective, our results may be viewed as a justification for the $U$-hat method in the special and unrealistic case of linear outcome model (2).
-In general, however, our method is fundamentally different from the U-hat method.
-The U-hat method relies on regression adjustment, which assumes a specified outcome-confounder relationship.
-If confounders $U$ were observed, then, at least in principle, this relationship could be estimated from the data.
-But if $U$ is unobserved, then specifying the functional form of $g$ depends entirely on subject-matter knowledge, which is rarely available.
-In Section LABEL:sec:Uhat of the supplementary material, we explain in detail why the $U$-hat method fails under the general model $Y=f(X;\beta)+g(U)+\epsilon_{y}$.
-In contrast, our approach constructs instrumental variables, which are agnostic to the form of unmeasured confounding even if the causal relationship is non-linear.
+From this perspective, our results may be viewed as a justification for the $U$-hat method in the special and unrealistic case of the linear outcome model (2).
+In general, our approach differs fundamentally from the U-hat methods.
+The U-hat1 method yields consistent estimators of $\beta$ only when the treatment–outcome relationship is linear.
+As shown in Section LABEL:sec:Uhat of the supplementary material, U-hat1 fails under the more general model
+$Y=f(X;\beta)+g(U)+\epsilon_{y},$
+where $f(X;\beta)$ may be nonlinear.
+In particular, we derive a necessary condition (Equation (LABEL:eqn:identifiabilityequation_supp)) that must hold for U-hat1 to consistently estimate $\beta$, but this condition is typically violated in nonlinear outcome models.
+The U-hat2 method, on the other hand, relies on modeling the latent confounder–outcome relationship, introducing additional assumptions that are implausible in practice and do not resolve the underlying identification challenge.
+In contrast, our approach constructs instrumental variables, which are agnostic to the form of unmeasured confounding, even if the causal relationship is nonlinear.
 Because of this, we believe it is a more robust approach in this setting.
 To our knowledge, we are the first to apply the IV framework to this problem, offering a method that accommodates nonlinear causal effects and arbitrary dependence on unmeasured confounders $U$.
-See Section LABEL:sec:Uhat in the supplementary materials for more discussions.
+See Section LABEL:sec:Uhat of the supplementary material for further discussion and a comparison between the proposed method and the U-hat methods.
 (a) Non-linear setting 1.
-(b) Non linear setting 2.
+(b) Non-linear setting 2.
 […]
@@
 Under the plurality rule A4, we conclude that the causal effects are identifiable as $\widehat{q}+\widehat{s}=8\ll p=2819$.
-Specifically, our analysis suggests that increasing the gene expression levels of $Igfbp2$, $Rab27a$, $Dct$, $Ankhd1$, and $Gck$ by one standard deviation leads to changes of $-1.98$, $1.88$, $1.43$, $-1.33$, and $1.17$ grams in mouse body weight, respectively.
+Specifically, our analysis suggests that increasing the gene expression levels of $Igfbp2$, $Rab27a$, DctDct, $Ankhd1$, and $Gck$ by one standard deviation leads to changes of 1.98([2.69,1.27])-1.98([-2.69,-1.27]), $1.88([1.26,2.51])$, $1.43([0.86,2.02])$, 1.33([1.90,0.77])-1.33([-1.90,-0.77]), and $1.17([0.69,1.66])$ grams in mouse body weight, respectively.
+The empirical confidence intervals were constructed using the ivreg function; see Section LABEL:sec:CI_sup for more details.
 We compared our approach with six methods: (i) the Lasso method; (ii) the two-stage regularization (2SR) method (Lin et al., 2015), which leverages SNPs as high-dimensional instrumental variables to estimate the causal effect; (iii) the auxiliary variable method (Miao et al., 2023), which focuses on the 17 genes identified by Lin et al.
@@
 (2015) as auxiliary variables; (iv) the null variable method (Miao et al., 2023), which assumes that more than half of the 17 genes have zero effects on mouse body weight; (v) the Trim method (Ćevid et al., 2020); and (vi) the IV-Lasso method, as detailed in Section 5.
-Detailed results for these comparison methods are included in Supplementary Material Section LABEL:sec:comparison.
-The number of active genes found by these methods, defined as genes that have non-zero effects on mouse body weight, is 87, 17, 4, 2, 4, and 14, respectively.
+Detailed results for these comparison methods are included in Section LABEL:sec:comparison of the supplementary material.
+The number of active genes found by these methods, defined as genes with non-zero effects on mouse body weight, is 87, 17, 4, 2, 4, and 14, respectively.
 Consistent with the simulation results, the Lasso method identifies many more active exposures than our method, and its selected set includes all genes identified by our approach.
@@
 Our key assumption is a sparse causation assumption, which in many contexts serves as an appealing alternative to the widely adopted sparse association assumption.
-We develop a synthetic instrument approach to identify and estimate causal effects without the need for collecting additional exogenous information, such as instrumental variables.
+We develop a synthetic instrument approach to identify and estimate causal effects without the need to collect additional exogenous information, such as instrumental variables.
 Our estimation procedure can be formulated as an $\ell_{0}$-optimization problem and can therefore be solved efficiently using off-the-shelf packages.
@@
 We have focused on a linear treatment model (1).
-In a general nonlinear treatment model, where the relationship between treatment $X$ and the unmeasured factor $U$ is nonlinear, nonlinear factor analysis could be used to fit $X=m(U;\alpha)+\epsilon$.
+In a general nonlinear treatment model, where the relationship between treatment $X$ and the unmeasured factor $U$ is nonlinear, nonlinear factor analysis could be used to fit $X=m(U;\Lambda)+\epsilon$.
 However, identifying a function $h(\cdot)$ such that $h(X)$ is independent of $U$ remains a significant challenge.
@@
 We thank Xin Bing, Zhichao Jiang, Wang Miao, Thomas Richardson, James Robins, Dominik Rothenhäusler, and Xiaochuan Shi for their helpful discussions and constructive comments.
-We also extend our gratitude to the Editor, Associate Editor, and anonymous referees for their valuable and thoughtful input, which has significantly enhanced the quality of this manuscript.
+We also extend our gratitude to the Editor, the Associate Editor, and the anonymous referees for their valuable and thoughtful input, which has significantly improved the quality of this manuscript.
 Supplementary material